### v1

# MGDepth: Motion-Guided Cost Volume For Self-Supervised Monocular Depth In Dynamic Scenarios

Abstract
Despite advancements in self-supervised monocular depth estimation, challenges persist in dynamic scenarios due to the dependence on assumptions about a static world. In this paper, we present MGDepth, a Motion-Guided Cost Volume Depth Net, to achieve precise depth estimation for both dynamic objects and static backgrounds, all while maintaining computational efficiency. To tackle the challenges posed by dynamic content, we incorporate optical flow and coarse monocular depth to create a novel static reference frame. This frame is then utilized to build a motion-guided cost volume in collaboration with the target frame. Additionally, to enhance the accuracy and resilience of the network structure, we introduce an attention-based depth net architecture to effectively integrate information from feature maps with varying resolutions. Compared to methods with similar computational costs, MGDepth achieves a significant reduction of approximately seven percent in root-mean-square error for self-supervised monocular depth estimation on the KITTI-2015 dataset.
keywords:
Unsupervised Learning, Depth Estimation, Ego-motion Estimation, Dynamic Scenario, Attention Mechanism.
[…]
[label1]organization=University Of Oxford, city=Oxford, postcode=OX1 2JD, country=United Kingdom
1 Introduction
The role of vision-based depth estimation (VDE) has become increasingly important in computer vision due to its ability to understand the 3D geometry of a scene based on the 2D observation, which serves as the foundation for various high-level 3D tasks such as scene reconstruction niemeyer2022regnerf ; zhou2023dynpoint ; li2021hierarchical , object detection wang2021salient and navigation de2021deep . Moreover, VDE has enabled state-of-the-art applications ranging from autonomous driving you2019pseudo ; wang2019pseudo ; miao2023ds to augmented reality newcombe2011dtam ; mieloch2022overview ; noraky2019low .
Figure 1: Qualitative comparison on KITTI-2015. (a) shows the RGB image of the current frame. (b) displays the estimated depth generated by Monodepth2 trained with monocular mode. (c) presents the estimated depth of Monodepth2 trained with stereo mode. (d) demonstrates the estimated depth of MGDepth in monocular mode. Differences are masked out by circles.
Recently, self-supervised depth estimation has emerged as a viable approach for training depth estimation methods, aiming to alleviate the dependency on extensive training data and reduce high computational demands. These methods learn depth maps from either monocular images bian2021unsupervised or stereo image pairs yao2018mvsnet . Despite the significant advancements made in self-supervised monocular vision-based depth estimation, a notable performance gap persists when comparing self-supervised monocular VDEs to self-supervised stereo VDEs. The disparity in performance can be mainly ascribed to the capability of stereo methods to utilize multiple views for constructing a feature volume, thereby incorporating a greater amount of 3D camera frustum information. While multi-frame monocular VDEs, as presented in watson2021temporal , have the ability to construct a feature volume based on adjacent frames, the presence of dynamic elements in these adjacent frames can potentially disrupt the construction of the feature volume.
Taking the aforementioned concern into account, this paper strives to enhance the performance of multi-frame monocular VDEs by integrating motion information during the inference process and adeptly managing the influence of dynamic objects in the construction of the feature volume. Experimental results indicate that the proposed multi-frame monocular VDEs can surpass even previous monocular VDEs that directly incorporate stereo information during the training process, as demonstrated in the examples provided in Fig. 1.
Specifically, in light of recent advancements in optical flow estimation and its successful application in motion detection and estimation tasks, we propose MGDepth. MGDepth is a self-supervised monocular vision-based depth estimation (VDE) system that leverages motion information through a motion-guided cost volume constructed with an attention mechanism. The attention mechanism is chosen due to its demonstrated outstanding performance in representation learning and effective fusion of diverse information. Our main contributions are summarized as follows:
  • 1.
    We utilize estimated optical flow alongside prior depth information to generate a new static reference frame. This reference frame effectively neutralizes the influence of dynamic elements within the original frame.
  • 2.
    By incorporating the new static reference frame, the target frame, and the initial reference frame, we construct a novel motion-guided volume that captures the dynamics of moving objects.
  • 3.
    Leveraging the High-Resolution network (HRNet), we introduce an innovative depth estimation architecture that employs attention mechanisms. This enables the integration of features with varying levels of detail, resulting in precise pixel-wise dense predictions.
  • 4.
    Our proposed model outperforms existing single and multi-frame methods on the KITTI, Cityscapes, and Odometry datasets. Additionally, our model can be trained efficiently using only a single NVIDIA A10 graphics card within a reasonable timeframe.
The structure of this paper unfolds as subsequent sections detail various aspects. Section II offers a concise overview of related studies concerning monocular depth estimation techniques. Section III delves into the intricacies of the proposed MGDepth. Findings from experiments and a thorough model analysis are outlined in Section IV. Section V wraps up with the conclusions drawn from the presented research.
2 Related Works
2.1 Monocular Depth Estimation
Self-supervised techniques for monocular depth prediction have gained notable traction due to their independence from labeled data and their versatility across various scenarios. Broadly, monocular VDEs fall into two categories: one relies solely on the present frame for depth estimation, while the other employs multiple neighboring frames to achieve depth prediction.
2.1.1 Single-frame Monocular Depth Estimation
In the initial category, Monodepth zhou2017unsupervised emerges as a leading benchmark, showcasing reliable results. The model features two networks: one focused on pose estimation and the other on depth estimation. During the optimization process, these networks synergize by utilizing the warping connection between depth and image transformations to construct a photometric loss function. Building upon this framework, Monodepth2 godard2019digging introduces a minimal projection error approach to tackle occlusion and minimize visual artifacts through full-resolution multiscale sampling. In addition to employing traditional photometric loss, FeatDepth shu2020feature introduces a feature-space-based metric loss to more effectively regulate the warping relationship between target and source frames. Moreover, DevNet zhou2022devnet enhances its performance by integrating 3D geometric consistency among neighboring frames.
2.1.2 Multi-frame Monocular Depth Estimation
The premise of the second category rests on the notion that integrating temporal information during inference—by employing multiple neighboring frames as inputs—can enhance the accuracy of the final depth estimation. Initially, this is accomplished by employing test-time refinement techniques casser2019depth ; chen2019self ; mccraith2020monocular ; kuznietsov2021comoda , along with recurrent neural networks (RNNs) as evidenced in studies like patil2020don ; wang2019recurrent ; zhang2019exploiting . The test-time refinement method adopts a monocular strategy to leverage temporal data during testing, whereas the recurrent neural network integrates with a monocular depth estimation network to analyze continuous frame sequences. Nonetheless, models utilizing recurrent neural networks often entail high computational costs and lack a distinct geometric inference approach. Recently, Manydepth watson2021temporal and MonoRec wimbauer2021monorec have made notable advancements in performance and real-time efficiency by incorporating cost volumes from stereo-matching tasks for geometric-based reasoning gu2020cascade . These models rely on a photometric loss function, where temporally neighboring frames are mapped onto the current image plane using predetermined depth bins. Within the cost volume framework, the inferred depth with the minimal value corresponds closely to the actual depth.
Nevertheless, these approaches are grounded in static assumptions regarding scenarios and struggle with dynamic foreground elements. To tackle this limitation, we introduce MGDepth, a technique adept at managing dynamic foreground by integrating temporal data into the cost volume and implementing a motion-guided photometric loss function.
Figure 2: Illustration for the structure of MGDepth. During Stage 1, target and reference frame $I_{t}$ and $I_{r}$ are processed Flow Net, Pose Net, and Depth Net to generate optical flow $f_{r\leftarrow t}$, transformation matrix $[R|\bm{t}]_{rt}$, and coarse depth $D^{c}_{t}$. During Stage 2, the outputs are used to generate the motion-guided cost volume $\mathcal{V}_{m}$. Finally, the motion-guided cost volume $\mathcal{V}_{m}$ and the target frame $I_{t}$ are used by Attention-Based Depth Net to produce the refined depth $D_{t}$.
2.2 Self-supervised VDE for Dynamic Objects
Due to the distinctive characteristics of dynamic objects, researchers have pursued differential treatment of these objects during both the training and inference processes in self-supervised vision-based depth estimation (VDE) methods. In klingner2020self , dynamic objects are deliberately excluded from consideration to prevent their influence on the optimization process and maintain the accuracy of the depth estimation. In feng2022disentangling ; ranftl2016dense , dynamic objects are initially segmented and subsequently treated differently in the photometric loss function. This approach allows for the incorporation of dynamic objects into the depth learning process, enabling their meaningful contribution rather than direct exclusion. Moreover, in lee2021learning ; lee2021attentive ; li2021unsupervised , an effort is made to predict object-level motion information and leverage it to provide enhanced constraints for self-supervised depth learning. However, these methods either suffer from high computational demands or exhibit limited performance improvement. Our proposed method, MGDepth, addresses these limitations by constructing a motion-guided cost volume that integrates predicted optical flow information through an attention mechanism.
3 Method
3.1 Overview
The primary structure of MGDepth is based on previous self-supervised monocular depth estimation algorithms, such as Monodepth godard2017unsupervised and Manydepth watson2021temporal , with the goal of estimating depth from monocular video recorded by an unknown camera. However, existing methods face difficulties in accurately estimating depth for moving objects lee2021learning ; li2021unsupervised . We propose the following innovative strategies to achieve accurate monocular depth estimation in complex dynamic scenarios. (1) We make use of the optical flow information and prior depth information to construct a new static reference frame. (2) Based on the new static reference frame, the original reference frame, and the new target frame, we construct a motion-guided cost volume. (3) We introduce an attention-based depth net structure, specifically designed to proficiently amalgamate feature maps with varying resolutions derived from a high-resolution network.
3.2 Framework
Our objective is to estimate the depth map ${D_{t}}$ and the rigid transformation ${[R|\bm{t}]_{rt}}$ between the target frame ${I_{t}}$ and a reference frame ${I_{r}}$. To create the constraint for self-supervision, we can first reconstruct target frame ${I^{\prime}_{t}}$ in the following manner:
$I^{\prime}_{t}=\phi(I_{r},D_{t},K_{t},K_{r},[R|\bm{t}]_{rt}).$ (1)
Here, ${K_{t},K_{r}}$ represents the intrinsic matrices of the target and reference frames, and $\phi(.)$ denotes the projection process based on homography warping. The ultimate photometric loss function, denoted as $\mathcal{L}_{p}$, is computed as the discrepancy between the transformed image ${I^{\prime}_{t}}$ and the reference image ${I_{t}}$. This loss function serves as the optimization objective for the proposed neural network.
The structure of MGDpth is demonstrated in Fig. 2. At the outset, MGDpth employs a pretrained optical flow network denoted by $\theta_{f}$ and a pose network denoted by $\theta_{p}$ to predict the optical flow $f_{r\leftarrow t}$ and the transformation matrix $[R|\bm{t}]_{rt}$ from the target frame to reference frame. Concurrently, a prior depth network denoted by $\theta_{cd}$ is utilized to estimate the coarse depth map $D_{t}^{c}$ for the target frame. Subsequently, the target frame denoted as $I_{t}$, the reference frame represented by $I_{r}$, the coarse depth map denoted as $D_{t}$, the transformation matrix $[R|\bm{t}]_{rt}$, and the optical flow $f_{r\leftarrow t}$ are utilized collaboratively to construct a motion cost volume denoted as $\mathcal{V}_{m}$. The motion-guided cost volume $\mathcal{V}_{m}$ and the target frame $I_{t}$ are used in unison as inputs to the attention-based depth network $\theta_{ad}$ with the objective of predicting the final depth map $D_{t}$. Subsequently, this predicted depth map $D_{t}$ is utilized in the construction of the ultimate motion-guided photometric loss denoted as $\mathcal{L}_{mp}$.
Figure 3: Relationship between optical flow and depth in the dynamic scenario. This figure demonstrates that in a dynamic scenario, there is a discrepancy between the static optical flow and the real optical flow. $p_{t}$,
3.3 New Static Reference Frame
To address the depth estimation of moving objects with monocular video, we leverage the estimation of flow $f_{r\leftarrow t}$, transformation matrix $[R|\bm{t}]_{rt}$, and coarse depth $D^{c}_{t}$. Using the warping relationship between $I_{t}$ and $I_{r}$, we can calculate a depth-based flow (static flow) $f^{\prime}_{r\leftarrow t}$ according to the following expression:
$f^{\prime}_{r\leftarrow t}=\frac{1}{D_{r}}(K_{r})(R_{rt}(K_{t})^{-1}D^{c}_{t}p_{t}+\bm{t}_{rt})-p_{t},$ (2)
where $p_{t}$ are pixels in the frame $I_{t}$; $K_{r}\&K_{t}$ are intrinsic matrix for frames $I_{r}\&I_{t}$. As illustrated in Fig. 3, in static scenarios, the static flow $f^{\prime}_{r\leftarrow t}$ aligns with the real optical flow $f_{r\leftarrow t}$. In scenarios involving moving objects, the real optical flow $f_{r\leftarrow t}$ can be decomposed into static optical flow $f^{\prime}_{r\leftarrow t}$ and dynamic optical flow $f^{d}_{r\leftarrow t}$. Then the dynamic flow can be calculated as $f^{d}_{r\leftarrow t}=f_{r\leftarrow t}-f^{\prime}_{r\leftarrow t}$. The motion mask could be generated as follow:
$\mathcal{M}_{m}^{t}=||f_{r\leftarrow t}-f^{\prime}_{r\leftarrow t}||^{2}_{2}>\epsilon,$ (3)
where $\epsilon$ is the threshold for distinguishing moving parts. The motion pattern observed in the reference frame can be utilized to create a corresponding mask in the target frame. This is achieved through the generation of a static motion mask $\mathcal{M}_{ms}^{r}$, utilizing the static optical flow $f^{\prime}_{r\leftarrow t}$. Simultaneously, the dynamic motion mask $\mathcal{M}_{md}^{r}$ is derived with the assistance of $f_{r\leftarrow t}$. The composite motion mask for the reference frame is obtained by combining these two masks, denoted as $\mathcal{M}_{m}^{r}=\mathcal{M}_{ms}^{r}\cup\mathcal{M}_{md}^{r}$. The rationale for avoiding the direct computation of mask $\mathcal{M}_{m}^{r}$ using $f^{\prime}_{t\leftarrow r}$ lies in the necessity of inferring the depth of the target frame $D_{r}$. This procedure would subsequently lead to a rise in computational demands for both the training and inference processes.
With $\mathcal{M}_{m}$, it becomes possible to reconstruct a new static reference image $\hat{I_{r}}$ by eliminating the impact of dynamic objects. This process can be formulated as follows:
$\hat{I_{r}}=\begin{cases}I_{r}&p_{i}\notin\mathcal{M}_{m};\\ \phi^{\prime}(I_{t},D_{t},K_{t},K_{r},[R|\bm{t}]_{rt})&p_{i}\in\mathcal{M}_{m}.\\ \end{cases}$ (4)
It is important to note that $\phi^{\prime}(.)$ differs from $\phi(.)$, as $\phi(.)$ is utilized to generate current frame information via sampling based on pixel correspondence between the current frame and the adjacent frame, whereas $\phi^{\prime}(.)$ generates adjacent frame by shifting forward pixels of the current frame. Please note that $\hat{I_{r}}$ is detached to eliminate its impact on the gradient calculation of backpropagation process.
3.4 Motion-Guided Cost Volume
Figure 4: Illustration for the construction of Final Volume. Feature maps of the target, reference, and new static reference frames $\mathcal{F}_{t}$, ${\mathcal{F}_{r}}$ and $\hat{\mathcal{F}_{r}}$ are used to generate motion-guided cost volume $\mathcal{V}_{m}$ with the help of $\mathcal{D}$ and $[R|t]_{rt}$. This volume $\mathcal{V}_{m}$ is combined with target feature map $\mathcal{F}_{t}$ via a channel attention to form the Final Volume.
In constructing the motion-guided cost volume $\mathcal{V}_{m}$ for target frame $I_{t}$, a set of parallel planes that are perpendicular to the optic axis of $I_{t}$ is defined, based on the depth assumptions $\mathcal{D}=\{d_{k};k=1,...,M\}$, where $M$ represents the number of planes. The feature extractor $\theta_{fe}$ is utilized to generate the feature maps $\mathcal{F}_{t}$, $\mathcal{F}_{r}$ and $\hat{\mathcal{F}_{r}}$ of $I_{t}$, $I_{r}$ and $\hat{I_{r}}$, respectively. With the aid of $\mathcal{D}$ and $[R|\bm{t}]_{rt}$, a set of wrapped feature maps $\{\hat{\mathcal{F}_{t\leftarrow r}^{d_{k}}},\mathcal{F}_{t\leftarrow r}^{d_{k}};d_{k}\in\mathcal{D}\}$ are generated by warping $\hat{\mathcal{F}_{r}}$ and $\mathcal{F}_{r}$. Considering $\mathcal{V}_{m}=\{\mathcal{P}_{k},k=1,...,M\}$, it could be written as:
$\mathcal{P}_{k}=\alpha\frac{\sum_{i=1}^{N}|\hat{\mathcal{F}_{t\leftarrow r}^{d_{k}}}-\mathcal{F}_{t}|}{N}+(1-\alpha)\frac{\sum_{i=1}^{N}|\mathcal{F}_{t\leftarrow r}^{d_{k}}-\mathcal{F}_{t}|}{N}.$ (5)
Figure 5: Illustration for the construction of Final Volume. Feature maps of the target, reference, and new static reference frames $\mathcal{F}_{t}$, ${\mathcal{F}_{r}}$ and $\hat{\mathcal{F}_{r}}$ are used to generate motion-guided cost volume $\mathcal{V}_{m}$ with the help of $\mathcal{D}$ and $[R|t]_{rt}$. This volume $\mathcal{V}_{m}$ is combined with target feature map $\mathcal{F}_{t}$ via a channel attention to form the Final Volume.
Here, $N$ denotes the number of reference frames used in construction, while in our experiment, we typically use a single adjacent frame as a reference frame. $\alpha$ denotes the parameters used to balance the influence of new static reference frame $\hat{I_{r}}$. The main reason for keeping the first item in the motion-guided cost volume is that at the early stage of the training process, the prior depth is not accurate enough which makes the $f^{\prime}_{r\leftarrow t}$ noisy. With the progress of training, $\alpha$ will gradually decrease to improve the influence of the new static reference frame in the construction of $\mathcal{V}_{m}$.
Instead of directly taking the volume $\mathcal{V}_{m}$ into the later depth network, we propose to construct the final feature volume $\mathcal{V}_{f}$ by combining the $\mathcal{V}_{m}$ with the $\mathcal{F}_{t}$ though channel attention. The process of constructing the final feature volume $\mathcal{V}_{f}$ is presented in Fig. 5.
Figure 6: Illustration for the structure of Attention-based Depth Net. To effectively leverage feature maps of varying resolutions, we employ efficient channel attention to concatenate high-resolution feature maps with low-resolution feature maps and the low-resolution disparity map. This resultant feature map is subsequently utilized for disparity prediction within the current branch and is propagated to the subsequent branch.
3.5 Attention-Based Depth Network
The High-Resolution Network (HRNet) wang2020deep is well-regarded for its ability to preserve a high level of detail in input images. Although it is widely used in various computer vision tasks, its performance in Vision Depth Estimation (VDE) has not been extensively explored. To enhance the accuracy and efficiency of VDE, we integrate an efficient channel attention module into the design of the depth decoder network of the HRNet-based depth network.
The HRNet is composed of multiple branches, denoted as $B$, each generating $S=b,...,4$ features with resolutions of $(\frac{H}{2^{b-1}},\frac{W}{2^{b-1}})$. Following the framework of Monodepth2, we utilize all branches to produce multi-scale disparity maps denoted as $D^{b}$, and apply the same loss function to each branch.
However, instead of exclusively utilizing the feature map from the last stage of each branch for disparity map prediction, we leverage the efficient channel attention mechanism. This mechanism integrates feature maps from the current branch’s various stages, feature maps from deeper branches, and disparities from deeper branches. This fusion process can be expressed as $\mathcal{A}([f^{b}_{s}]_{s=b}^{s=4},[f^{b+1}_{s}]_{s=b+1}^{s=4},D^{b+1})$, as in Fig. 6 for the $b^{th}$ branch, to collectively inform the disparity estimation process.
3.6 Loss Function
To summarize, there are three loss functions used in updating MGDepth’s weights. The final loss function $\mathcal{L}$ could be written as:
$\mathcal{L}=\mathcal{L}_{p}+\mathcal{L}_{s}+\mathcal{L}_{c}.$ (6)
The photometric loss $\mathcal{L}_{p}$ consists of $L_{1}$ norm and SSIM regularization:
$\begin{split}\mathcal{L}_{p}&=0.15*\sum_{p}\mathcal{M}_{o}\odot|I_{t}-I^{\prime}_{t}|\\ &+0.85*\frac{1-\text{SSIM}(I_{t},I^{\prime}_{t})}{2},\end{split}$ (7)
where $\mathcal{M}_{o}$ is the auto mask introduced in godard2017unsupervised . $\mathcal{L}_{s}$ is the smooth loss, which could be written as:
$\mathcal{L}_{s}=|\partial_{x}D_{t}^{*}|e^{-|\partial_{x}I_{t}|}+|\partial_{y}D_{t}^{*}|e^{-|\partial_{y}I_{t}|},$ (8)
where $d_{t}^{:}*=d_{t}/\overline{d_{t}}$ and $\overline{d_{t}}$ is the average depth of target frame wang2018learning . And $\mathcal{L}_{c}$ is the consistency loss introduced by watson2021temporal , to preserve the consistency between prior monocular depth and final multi-frame depth, which could be written as:
$\mathcal{L}_{c}=\sum_{p}\mathcal{M}_{m}^{t}\odot|D_{t}^{c}-D_{t}|.$ (9)
3.7 Network Structure
3.7.1 Depth Prior Network
The architecture of the Depth Prior Network (DPN) encompasses a dual-component system consisting of a depth encoder and a depth decoder. The depth encoder is established employing the HRNet16 model wang2020deep , which has been pre-trained on the ImageNet dataset deng2009imagenet and functions as the foundational backbone. This encoder operates across four distinct branches, each corresponding to a distinct pyramidal scale. The feature representations within each of these scales are subsequently channeled to the depth decoder through a sequence of skip connections, notably influenced by the UNet framework ronneberger2015u .
The depth decoder is characterized by an attention-based architecture as originally introduced in our primary manuscript. This component is designed to facilitate the fusion of encoded features derived from the depth encoder.
3.7.2 Pose Network
PoseNet exclusively comprises an encoder architecture utilizing a ResNet18 model he2016deep that has been pre-trained on the ImageNet dataset deng2009imagenet . Nonetheless, its distinctive attribute lies in its capacity to generate a 6-degree-of-freedom camera ego-motion vector, including a 3d translational vector and a 3d rotational vector.
3.7.3 Depth Network
In contrast to the Depth Prior Network, the depth network has a very similar layout. The key distinction is that, rather than analyzing just one frame at a time, it often processes two consecutive frames together. Once the first stage’s feature extraction is complete, these paired feature maps are employed to create a ”cost volume” as outlined in the main text. This cost volume is then blended with the feature representation of the target frame, resulting in a comprehensive ”final volume.” This final volume plays a crucial role in the subsequent process of building the depth map.
3.7.4 Optical Flow Network
The low-level computer vision community has favored the task of estimating the pixel-wise correspondence (Optic Flow) between two frames due to its applicability across various scenarios, without the need for specialized prior knowledge. Our paper involves adapting the pre-trained Gmflow xu2022gmflow on synthetic datasets, including the Flying Thing, Monkaa, and Driving datasets from the Scene Flow Dataset MIFDB16 and the Sintel dataset butler2012naturalistic . Notably, Gmflow used in MGDepth was not trained on any dataset used in our depth and ego-motion estimation experiments. The rationale behind adopting Gmflow lies in its lightweight nature and rapid inference capabilities, attributed to the inclusion of a dilatation correlation layer. This selection was made based on the consideration that integrating Gmflow into the existing depth estimation framework would not impose substantial computational overhead.
Figure 7: Qualitative results on KITTI Dataset. The initial row displays the RGB images of target frames where depth information has been estimated. The subsequent rows, specifically the second, third, and fourth, showcase depth maps generated by Monodepth2 trained with stereo techniques, ManyDepth, and MGDepth, respectively.
3.8 Discussion And Thinking
MGDepth distinguishes itself from previous methods, such as feng2022disentangling , that utilize category information for segmentation and object tracking. Unlike those methods, MGDepth operates independently of any prior dataset knowledge. The optical flow module, pretrained solely on synthetic testing data, significantly enhances its ability to adapt to various scenarios, eliminating the need for extra labeling efforts.
Furthermore, unlike typical motion segmentation modules relying on optical flow yang2021learning that demand substantial computational resources, MGDepth takes a different approach. We directly leverage the optical flow model and integrate necessary information, resulting in reduced computational demands.
4 Experiment
The primary experiments were conducted on the KITTI-2015 geiger2012we and Cityscapes silberman2012indoor datasets. (1) KITTI-2015: We use the Eigen split eigen2015predicting for both training and testing our model. To maintain consistency with prior studies zhou2017unsupervised ; zhou2021r , we utilized the same data selection approach and excluded all static images from the training set. This resulted in 39,810 training images and 4,424 test images.
To ensure uniformity, we employed the camera matrix configuration described in godard2019digging , which applies a single intrinsic matrix to all images and sets the length focal to the average value.
(2) Cityscapes: To ensure consistency with previous investigations watson2021temporal ; yang2018lego , we selected 69,731 images from the monocular sequences from Cityscapes cordts2016cityscapes as our training dataset, which were preprocessed into triples using scripts from zhou2017unsupervised . Our approach does not take into account stereo pairs or semantics. Performance assessment was based on the 1,525 test images, and the provided SGM hirschmuller2007stereo disparity maps were used. Similar to KITTI, we restricted depths to a max of 80m.
T Method M S F W x H Abs Rel Sq Rel RMSE RMSE log $\delta<1.25$ $\delta<1.25^{2}$ $\delta<1.25^{3}$
KITTI 2015
Stereo
EPC++luo2019every 1024 x 320 0.128 0.935 5.011 0.209 0.831 0.945 0.979
FeatDepthshu2020feature 1024 x 320 0.099 0.697 4.427 0.184 0.889 0.963 0.982
R-MSFM6zhou2021r 1024 x 320 0.108 0.753 4.469 0.185 0.888 0.963 0.982
DFRzhan2018unsupervised 608 x 160 0.135 1.132 5.585 0.229 0.820 0.933 0.971
Monocular
Packnet-SFM guizilini20203d 1280 x 384 0.107 0.802 4.538 0.186 0.889 0.962 0.981
Monodepth2 godard2019digging 1024 x 320 0.115 0.882 4.701 0.190 0.879 0.961 0.982
Zhao zhao2020towards $\bullet$ 1024 x 320 0.113 0.704 4.581 0.184 0.871 0.961 0.984
Chen chen2022compacted $\bullet$ $\bullet$ 1024 x 320 0.112 0.859 4.798 0.193 0.879 0.960 0.981
FeatDepth shu2020feature 1024 x 320 0.104 0.729 4.481 0.179 0.893 0.965 0.984
Wang wang2020self $\bullet$ 1024 x 320 0.106 0.773 4.491 0.185 0.890 0.962 0.982
DevNet zhou2022devnet 1024 x 320 0.103 0.713 4.459 0.177 0.890 0.965 0.982
Ranjan ranjan2019competitive $\bullet$ 832 x 256 0.148 1.149 5.464 0.226 0.815 0.935 0.973
EPC++ luo2019every 832 x 256 0.141 1.029 5.350 0.216 0.816 0.941 0.976
Guizilini guizilini2020semantically $\bullet$ 640 x 192 0.102 0.698 4.381 0.178 0.896 0.964 0.984
Johnston johnston2020self 640 x 192 0.106 0.861 4.699 0.185 0.889 0.962 0.982
Monodepth2 godard2019digging 640 x 192 0.115 0.903 4.863 0.193 0.877 0.959 0.981
Packnet-SFM guizilini20203d 640 x 192 0.111 0.785 4.601 0.189 0.878 0.960 0.982
Patil patil2020don $\bullet$ $\bullet$ 640 x 192 0.111 0.821 4.650 0.187 0.883 0.961 0.982
Wang wang2020self $\bullet$ 640 x 192 0.106 0.799 4.662 0.187 0.889 0.961 0.982
Li li2023learning $\bullet$ $\bullet$ 640 x 192 0.102 0.703 4.348 0.175 0.895 0.966 0.984
ManyDepthwatson2021temporal $\bullet$ 640 x 192 0.098 0.770 4.459 0.176 0.900 0.965 0.983
Lee lee2023multi $\bullet$ 640 x 192 0.096 0.644 4.230 0.172 0.903 0.968 0.985
DynamicDepth feng2022disentangling $\bullet$ $\bullet$ 640 x 192 0.096 0.720 4.458 0.175 0.897 0.964 0.984
MGDepth $\bullet$ $\bullet$ 640 x 192 0.091 0.650 4.263 0.171 0.908 0.967 0.984
Cityscapes
Monocular
Lee lee2021attentive 832 x 256 0.116 1.213 6.695 0.186 0.852 0.951 0.982
InstaDM lee2021learning $\bullet$ 832 x 256 0.111 1.158 6.437 0.182 0.868 0.961 0.983
Pilzer pilzer2018unsupervised 512 x 256 0.240 4.264 8.049 0.334 0.710 0.871 0.937
Monodepth2 godard2019digging 416 x 128 0.129 1.569 6.876 0.187 0.849 0.957 0.983
Videos in the Wild gordon2019depth 416 x 128 0.127 1.330 6.960 0.195 0.830 0.947 0.981
Li li2021unsupervised 416 x 128 0.119 1.290 6.980 0.190 0.846 0.952 0.982
Struct2Depth casser2019depth $\bullet$ 416 x 128 0.151 2.492 7.024 0.202 0.826 0.937 0.972
ManyDepthwatson2021temporal $\bullet$ 416 x 128 0.114 1.193 6.223 0.170 0.875 0.967 0.989
DynamicDepth feng2022disentangling $\bullet$ $\bullet$ 416 x 128 0.103 1.000 5.867 0.157 0.895 0.974 0.991
MGDepth $\bullet$ $\bullet$ 416 x 128 0.097 0.792 5.827 0.154 0.903 0.975 0.993
Table 1: Self-supervised depth estimation results on the KITTI 2015 and Cityscapes. We evaluate the performance of algorithms utilizing either stereo or monocular information during the training phase. Metrics denoted by Color have a lower optimal value, while metrics denoted by Color have a higher optimal value. The upper portion shows results on the KITTI 2015 dataset and the lower portion shows results on the Cityscapes dataset. The optimal results in each subsection are denoted in bold. Our method outperforms all prior approaches in most metrics across all subsections, irrespective of whether the baselines utilize multiple frames during testing. M: Multi-frames. S: Motion Segmentation. F: Optical Flow.
(3) KITTI Odometry: For the purpose of odometry evaluation, we subject our system to the official KITTI odometry split, encompassing 11 driving sequences. These sequences are accompanied by ground truth odometry acquired through IMU and GPS readings. In line with prior research endeavors zhou2017unsupervised ; zhou2021r , our model’s training is conducted on sequences 00-08, while sequences 09-10 are reserved for testing.
The primary focus of this section is to evaluate the ability of MGDepth and other state-of-the-art methods to predict depth information and estimate ego-motion. The depth prediction results are evaluated using widely used metrics godard2019digging ; eigen2014depth .
The initial evaluation involves assessing MGDepth’s depth estimation performance on the KITTI-2015 and Cityscapes. Next, we evaluate its ability to estimate depth for dynamic objects on Cityscapes, utilizing the ground truth motion segmentation map. Following this, we conduct ablation studies to investigate the effectiveness of each module proposed in our paper. Finally, we present the results of MGDepth’s ego-motion estimation performance on the KITTI Odometry dataset.
4.1 Monocular Depth Estimation
KITTI-2015: In this experiment, the ResNet18 he2016deep architecture was utilized as the backbone for the pose network. The HRNet16 was used as the backbone for coarse depth network and attention-based depth network. During the training phase, our MGDepth receives input images with a resolution of $192\times 640$. For testing, we set the minimum and maximum depth values to 0.1m and 80m, respectively, following the recommendations of godard2019digging ; jin2001real . In our experiment, the scale factor is computed using the median value of the ground-truth image, employing the same approach as in zhou2017unsupervised . The upper section of Tab. 1 presents a comparison between the performance of MGDepth and state-of-the-art algorithms. Overall, MGDepth exhibits superior performance compared to other methods, irrespective of whether monocular or multiview images are used. In particular, MGDepth demonstrates a significant performance advantage over Manydepth, which also employs multiview images and a cost volume structure to estimate depth. MGDepth outperforms Manydepth by a substantial margin of $7.2\%$ in terms of Abs Rel. In comparison to Dynamicdepth, which employs semantic motion segmentation maps to leverage high-level computer vision information for precise depth estimation, MGDepth achieves a superior performance of $5.3\%$ in terms of the Abs Rel, without the inclusion of any domain-specific information. Quantitative results are shown in Fig. 7.
Cityscapes: In contrast to the KITTI dataset, the Cityscapes dataset features a higher percentage of dynamic objects. During the training phase on the Cityscapes dataset, the input resolution of MGDepth is set to $128\times 416$. The lower section of Tab. 1 indicates a substantial performance advantage of MGDepth over its competing methods. Remarkably, MGDepth exhibits superior performance over ManyDepth, with a margin of approximately $15.0\%$ in the absolute relative error metric. Through the visualization of error maps generated by computing the differences between predicted depth maps and their corresponding ground truth depth maps, Fig. 8 facilitates a qualitative analysis of depth predictions. These observations indicate that MGDepth offers enhanced accuracy for depth estimation of both dynamic foreground and static background. Additionally, the incorporation of a pre-trained optical flow estimation model with low computational requirements has resulted in a minimal increase in parameter size and running time for MGDepth. Quantitative results are shown in Fig. 8.
Type Method M WxH Abs Rel Sq Rel RMSE RMSE log $\delta<1.25$ $\delta<1.25^{2}$ $\delta<1.25^{3}$
S InstaDM lee2021learning 832 x 256 0.139 1.698 5.760 0.181 0.859 0.959 0.982
w/o S Monodepth2 godard2019digging 416 x 128 0.159 1.937 6.363 0.201 0.816 0.950 0.981
ManyDepth watson2021temporal $\bullet$ 416 x 128 0.169 2.175 6.634 0.218 0.789 0.921 0.969
MGDepth $\bullet$ 416 x 128 0.123 1.260 4.519 0.144 0.869 0.971 0.989
Table 2: Results of Dynamic Foreground on Cityscapes. We evaluate depth prediction errors for dynamic objects such as vehicles, people, and bikes on the Cityscapes dataset silberman2012indoor . The best results are highlighted in bold.
Figure 8: Qualitative results on Cityscapes (Error Map). The first row exhibits the RGB image of target frames for which depth information is estimated. The second and third rows illustrate error maps depicting the absolute relative error concerning the ground truth, where the degree of accuracy is represented by white (good) to red (poor) color gradations. The mask proposed in watson2021temporal is employed during drawing.
4.2 Depth For Dynamic Foreground
Tab. 2 showcases the depth errors within dynamic object regions by utilizing the ground truth motion segmentation maps provided in lee2021learning , which corresponds to the qualitative results presented in Fig. 8. Based on these results, it can be inferred that self-supervised monocular depth estimation methods from previous studies struggle to accurately estimate depth in dynamic scenes. This is likely due to their reliance on photometric loss, which assumes that the majority of foreground objects are static. In dynamic scenes, the photometric loss assumption is often violated, leading to difficulty in providing accurate depth estimates. However, MGDepth presents an effective solution to this issue by handling dynamic objects. Moreover, the motion-guided ego-motion loss function used in MGDepth enables accurate pose estimation, which aids in constructing cost volumes similar to stereo depth estimation. As a result, MGDepth achieves superior performance.
Figure 9: Trajectory visualization on Seq. 09. and Seq. 10. The ground truth trajectories are represented by red lines. Trajectories generated by the FeatDepth shu2020feature results are indicated by green lines, while our method’s trajectories are depicted in blue lines.
Method Tr R Tr R
DFR zhan2018unsupervised 11.93 3.91 12.45 3.46
MonoDepth2 godard2019digging 10.85 2.86 11.60 5.72
NeuralBundler li2019pose 8.10 2.81 12.90 3.71
SC-SfMlearner zhou2017unsupervised 8.24 2.19 10.70 4.58
FeatDepth shu2020feature 8.73$\dagger$ 2.11 10.65$\dagger$ 4.91
ManyDepthwatson2021temporal 8.08 1.97 9.86 3.42
MGDepth 7.01 1.76 7.29 2.65
Table 3: Visual odometry results on Seq. 9 and Seq. 10 of the KITTI odometry dataset. The average drift in root mean square error for both translation (Tr) and rotation (R) is reported. $\dagger$ denotes the method that outperforms the results reported in the original paper.
4.3 Odometry Estimation
To evaluate odometry estimation results, we follow the split used in zhou2017unsupervised ; zhan2018unsupervised . Specifically, we trained our model using Seq. 00-08 from the KITTI odometry dataset and test methods on Seq. 09-10. Tab. 3 shows the results, where the translational and rotational motion were evaluated by using the root mean square error (RMSE). MGDepth, which uses less computational resources, outperforms other learning-based techniques that use more computational resources, such as FeatDepth. The results indicate that MGDepth achieves a significant reduction in RMSE by approximately $24.1\%$ for translational motion and $22.5\%$ for rotational motion, as compared to ManyDepth, on Seq. 10. In support of this result, we present a visual depiction of the trajectory of both methods in Fig. 9, which shows that MGDepth’s trajectory exhibits considerably less drift than that of FeatDepth. The superior performance of MGDepth in precise pose estimation can be attributed to its use of both depth-based photometric loss and flow-based ego-motion loss, as well as the implementation of a motion-guided mask that effectively filters out dynamic foreground outliers. This combination of techniques enables MGDepth to achieve more reliable pose estimates compared to others.
Ablation Abs Rel Sq Rel RMSE
Final Prior Depth 0.108 0.753 4.563
w/o Motion-Guided Volume 0.094 0.676 4.289
w/o Attention for Final Volume 0.093 0.648 4.338
w Normal HRNet 0.096 0.746 4.349
MGDepth 0.091 0.650 4.263
Table 4: Ablation study results of depth accuracy on the KITTI 2015. The proposed MGDepth model was evaluated by modifying its components.
4.4 Abaltion Study
We conduct ablation studies on the Cityscapes dataset to demonstrate the effectiveness of our proposed modules and validate our design choices. The results are presented in Tab. 4, where the experiments were conducted under various settings, including Prior Depth Net of MGDepth, MGDepth without Motion-Guided Volume, MGDepth without Attention for Final Volume, and MGDepth with Normal HRNet. Tab. 4 indicates that the addition of Attention-Based HRNet has a significant impact on the performance of MGDepth. And the introduction of Motion-Guided Volume and the Attention for Final Volume could further improve the performance of MGDepth.
5 Conclusion
We present MGDepth, a self-supervised multi-frame monocular depth prediction model leveraging optical flow-depth geometry for enhanced depth estimation. Our approach involves novel components such as a static reference frame, motion-guided cost volume, and an attention-based high-resolution neural network. MGDepth achieves accurate depth estimation for dynamic foreground and static background on KITTI-2015 and Cityscapes datasets. Notably, our results are reproducible using a single NVIDIA A10 GPU. MGDepth seamlessly integrates with existing self-supervised monocular depth estimation algorithms, serving as a refined coarse depth network. Moreover, our proposed modules augment accuracy by incorporating dynamic information.
6 Acknowledgements
Our research is supported by Amazon Web Services in the Oxford-Singapore Human-Machine Collaboration Programme and by the ACE-OPS project (EP/S030832/1).
7 Declaration
During the preparation of this work the authors used ChatGPT 3.5 in order to improve language and readability. After using this tool/service, the author(s) reviewed and edited the content as needed and take(s) full responsibility for the content of the publication.

### v2

@@
-MGDepth: Motion-Guided Cost Volume For Self-Supervised Monocular Depth In Dynamic Scenarios
+Manydepth2: Motion-Aware Self-Supervised Monocular Depth Estimation in Dynamic Scenes
 Abstract
 Despite advancements in self-supervised monocular depth estimation, challenges persist in dynamic scenarios due to the dependence on assumptions about a static world.
-In this paper, we present MGDepth, a Motion-Guided Cost Volume Depth Net, to achieve precise depth estimation for both dynamic objects and static backgrounds, all while maintaining computational efficiency.
+In this paper, we present Manydepth2, a Motion-Guided Cost Volume Depth Net, to achieve precise depth estimation for both dynamic objects and static backgrounds, all while maintaining computational efficiency.
 To tackle the challenges posed by dynamic content, we incorporate optical flow and coarse monocular depth to create a novel static reference frame.
@@
 Additionally, to enhance the accuracy and resilience of the network structure, we introduce an attention-based depth net architecture to effectively integrate information from feature maps with varying resolutions.
-Compared to methods with similar computational costs, MGDepth achieves a significant reduction of approximately seven percent in root-mean-square error for self-supervised monocular depth estimation on the KITTI-2015 dataset.
-keywords:
-Unsupervised Learning, Depth Estimation, Ego-motion Estimation, Dynamic Scenario, Attention Mechanism.
-[…]
-[label1]organization=University Of Oxford, city=Oxford, postcode=OX1 2JD, country=United Kingdom
-1 Introduction
-The role of vision-based depth estimation (VDE) has become increasingly important in computer vision due to its ability to understand the 3D geometry of a scene based on the 2D observation, which serves as the foundation for various high-level 3D tasks such as scene reconstruction niemeyer2022regnerf ; zhou2023dynpoint ; li2021hierarchical , object detection wang2021salient and navigation de2021deep .
-Moreover, VDE has enabled state-of-the-art applications ranging from autonomous driving you2019pseudo ; wang2019pseudo ; miao2023ds to augmented reality newcombe2011dtam ; mieloch2022overview ; noraky2019low .
-Figure 1: Qualitative comparison on KITTI-2015.
-(a) shows the RGB image of the current frame.
-(b) displays the estimated depth generated by Monodepth2 trained with monocular mode.
-(c) presents the estimated depth of Monodepth2 trained with stereo mode.
-(d) demonstrates the estimated depth of MGDepth in monocular mode.
-Differences are masked out by circles.
+Compared to methods with similar computational costs, Manydepth2 achieves a significant reduction of approximately five percent in root-mean-square error for self-supervised monocular depth estimation on the KITTI-2015 dataset.
+The code could found here .
+Index Terms:
+Monocular Depth Estimation, Self-Supervised, Optical Flow.
+11footnotetext: KZ, JB, JZ, QX, NT and AM are with the University of Oxford. Contact Email: zhouk777@mit.edu
+I Introduction
+The role of vision-based depth estimation (VDE) has become increasingly important in computer vision due to its ability to understand the 3D geometry of a scene based on 2D observation, which serves as the foundation for various high-level 3D tasks, such as scene reconstruction [1], object detection [2] and navigation [3].
+Moreover, VDE has enabled state-of-the-art applications ranging from autonomous driving [4] to augmented reality [4].
 Recently, self-supervised depth estimation has emerged as a viable approach for training depth estimation methods, aiming to alleviate the dependency on extensive training data and reduce high computational demands.
-These methods learn depth maps from either monocular images bian2021unsupervised or stereo image pairs yao2018mvsnet .
+These methods learn depth maps from either monocular images [5] or stereo image pairs [6].
 Despite the significant advancements made in self-supervised monocular vision-based depth estimation, a notable performance gap persists when comparing self-supervised monocular VDEs to self-supervised stereo VDEs.
 The disparity in performance can be mainly ascribed to the capability of stereo methods to utilize multiple views for constructing a feature volume, thereby incorporating a greater amount of 3D camera frustum information.
-While multi-frame monocular VDEs, as presented in watson2021temporal , have the ability to construct a feature volume based on adjacent frames, the presence of dynamic elements in these adjacent frames can potentially disrupt the construction of the feature volume.
+While multi-frame monocular VDEs, as presented in [7], have the ability to construct a feature volume based on adjacent frames, the presence of dynamic elements in these adjacent frames can potentially disrupt the construction of the feature volume.
+Figure 1: Qualitative comparison on Cityscapes.
+The first row presents the RGB images.
+The second row shows the error maps of the depth estimates produced by ManyDepth.
+The third row displays the error maps of the depth estimates generated by Manydepth2.
 Taking the aforementioned concern into account, this paper strives to enhance the performance of multi-frame monocular VDEs by integrating motion information during the inference process and adeptly managing the influence of dynamic objects in the construction of the feature volume.
-Experimental results indicate that the proposed multi-frame monocular VDEs can surpass even previous monocular VDEs that directly incorporate stereo information during the training process, as demonstrated in the examples provided in Fig.
+Experimental results indicate that our method could effectively handle dynamic objects, as demonstrated in Fig.
 1.
-Specifically, in light of recent advancements in optical flow estimation and its successful application in motion detection and estimation tasks, we propose MGDepth.
-MGDepth is a self-supervised monocular vision-based depth estimation (VDE) system that leverages motion information through a motion-guided cost volume constructed with an attention mechanism.
+Specifically, in light of recent advancements in optical flow estimation and its successful application in motion detection and estimation tasks, we propose Manydepth2.
+Manydepth2 is a self-supervised monocular vision-based depth estimation system that leverages motion information through a motion-guided cost volume constructed with an attention mechanism.
 The attention mechanism is chosen due to its demonstrated outstanding performance in representation learning and effective fusion of diverse information.
 Our main contributions are summarized as follows:
-  • 1.
+Figure 2: Illustration for the structure of Manydepth2.
+During Stage 1, target and reference frame $I_{t}$ and $I_{r}$ are processed Flow Net, Pose Net, and Depth Net to generate optical flow $f_{r\leftarrow t}$, transformation matrix $[R|\bm{t}]_{rt}$, and coarse depth $D^{c}_{t}$.
+During Stage 2, the outputs are used to generate the motion-guided cost volume $\mathcal{V}_{m}$.
+Finally, the motion-guided cost volume $\mathcal{V}_{m}$ and the target frame $I_{t}$ are used by Attention-Based Depth Net to produce the refined depth $D_{t}$.
   • We utilize estimated optical flow alongside prior depth information to generate a new static reference frame.
 This reference frame effectively neutralizes the influence of dynamic elements within the original frame.
-  • 2.
   • By incorporating the new static reference frame, the target frame, and the initial reference frame, we construct a novel motion-guided volume that captures the dynamics of moving objects.
-  • 3.
   • Leveraging the High-Resolution network (HRNet), we introduce an innovative depth estimation architecture that employs attention mechanisms.
 This enables the integration of features with varying levels of detail, resulting in precise pixel-wise dense predictions.
-  • 4.
   • Our proposed model outperforms existing single and multi-frame methods on the KITTI, Cityscapes, and Odometry datasets.
 Additionally, our model can be trained efficiently using only a single NVIDIA A10 graphics card within a reasonable timeframe.
-The structure of this paper unfolds as subsequent sections detail various aspects.
-Section II offers a concise overview of related studies concerning monocular depth estimation techniques.
-Section III delves into the intricacies of the proposed MGDepth.
-Findings from experiments and a thorough model analysis are outlined in Section IV.
-Section V wraps up with the conclusions drawn from the presented research.
-2 Related Works
-2.1 Monocular Depth Estimation
+II Related Works
+II-A Monocular Depth Estimation
 Self-supervised techniques for monocular depth prediction have gained notable traction due to their independence from labeled data and their versatility across various scenarios.
-Broadly, monocular VDEs fall into two categories: one relies solely on the present frame for depth estimation, while the other employs multiple neighboring frames to achieve depth prediction.
-2.1.1 Single-frame Monocular Depth Estimation
-In the initial category, Monodepth zhou2017unsupervised emerges as a leading benchmark, showcasing reliable results.
-The model features two networks: one focused on pose estimation and the other on depth estimation.
-During the optimization process, these networks synergize by utilizing the warping connection between depth and image transformations to construct a photometric loss function.
-Building upon this framework, Monodepth2 godard2019digging introduces a minimal projection error approach to tackle occlusion and minimize visual artifacts through full-resolution multiscale sampling.
-In addition to employing traditional photometric loss, FeatDepth shu2020feature introduces a feature-space-based metric loss to more effectively regulate the warping relationship between target and source frames.
-Moreover, DevNet zhou2022devnet enhances its performance by integrating 3D geometric consistency among neighboring frames.
-2.1.2 Multi-frame Monocular Depth Estimation
-The premise of the second category rests on the notion that integrating temporal information during inference—by employing multiple neighboring frames as inputs—can enhance the accuracy of the final depth estimation.
-Initially, this is accomplished by employing test-time refinement techniques casser2019depth ; chen2019self ; mccraith2020monocular ; kuznietsov2021comoda , along with recurrent neural networks (RNNs) as evidenced in studies like patil2020don ; wang2019recurrent ; zhang2019exploiting .
+Broadly, monocular VDEs fall into two categories: one relies on the present frame for depth estimation, while the other employs multiple frames to achieve depth prediction.
+Single-frame Monocular Depth Estimation: Within the realm of depth estimation methodologies, the initial category has witnessed the emergence of notable frameworks that have set benchmarks in terms of reliability and performance.
+A prime example from this category is Monodepth, as highlighted in the study by [8].
+This model distinguishes itself with a bifurcated network architecture: one part is dedicated explicitly to pose estimation, while another part focuses on depth estimation tasks.
+A pivotal aspect of the Monodepth architecture lies in the collaborative synergy between these two networks.
+They operate by leveraging the warping connection intrinsic to depth and image transformations.
+This collaboration results in creating a photometric loss function, which is a key part of why it works so well.
+Further elevating the landscape of depth estimation methodologies, DevNet, in [9], takes a comprehensive approach by embedding 3D geometric consistency.
+Multi-frame Monocular Depth Estimation: The premise of the second category rests on the notion that integrating temporal information during inference—by employing multiple neighboring frames as inputs—can enhance the accuracy of the final depth estimation.
+Initially, this is accomplished by employing test-time refinement techniques [10], along with recurrent neural networks (RNNs) as evidenced in studies like [11].
 The test-time refinement method adopts a monocular strategy to leverage temporal data during testing, whereas the recurrent neural network integrates with a monocular depth estimation network to analyze continuous frame sequences.
 Nonetheless, models utilizing recurrent neural networks often entail high computational costs and lack a distinct geometric inference approach.
-Recently, Manydepth watson2021temporal and MonoRec wimbauer2021monorec have made notable advancements in performance and real-time efficiency by incorporating cost volumes from stereo-matching tasks for geometric-based reasoning gu2020cascade .
+Recently, Manydepth [7] and MonoRec [12] have made notable advancements in performance and real-time efficiency by incorporating cost volumes from stereo-matching tasks for geometric-based reasoning [13].
 These models rely on a photometric loss function, where temporally neighboring frames are mapped onto the current image plane using predetermined depth bins.
@@
 Nevertheless, these approaches are grounded in static assumptions regarding scenarios and struggle with dynamic foreground elements.
-To tackle this limitation, we introduce MGDepth, a technique adept at managing dynamic foreground by integrating temporal data into the cost volume and implementing a motion-guided photometric loss function.
-Figure 2: Illustration for the structure of MGDepth.
-During Stage 1, target and reference frame $I_{t}$ and $I_{r}$ are processed Flow Net, Pose Net, and Depth Net to generate optical flow $f_{r\leftarrow t}$, transformation matrix $[R|\bm{t}]_{rt}$, and coarse depth $D^{c}_{t}$.
-During Stage 2, the outputs are used to generate the motion-guided cost volume $\mathcal{V}_{m}$.
-Finally, the motion-guided cost volume $\mathcal{V}_{m}$ and the target frame $I_{t}$ are used by Attention-Based Depth Net to produce the refined depth $D_{t}$.
-2.2 Self-supervised VDE for Dynamic Objects
-Due to the distinctive characteristics of dynamic objects, researchers have pursued differential treatment of these objects during both the training and inference processes in self-supervised vision-based depth estimation (VDE) methods.
-In klingner2020self , dynamic objects are deliberately excluded from consideration to prevent their influence on the optimization process and maintain the accuracy of the depth estimation.
-In feng2022disentangling ; ranftl2016dense , dynamic objects are initially segmented and subsequently treated differently in the photometric loss function.
-This approach allows for the incorporation of dynamic objects into the depth learning process, enabling their meaningful contribution rather than direct exclusion.
-Moreover, in lee2021learning ; lee2021attentive ; li2021unsupervised , an effort is made to predict object-level motion information and leverage it to provide enhanced constraints for self-supervised depth learning.
-However, these methods either suffer from high computational demands or exhibit limited performance improvement.
-Our proposed method, MGDepth, addresses these limitations by constructing a motion-guided cost volume that integrates predicted optical flow information through an attention mechanism.
-3 Method
-3.1 Overview
-The primary structure of MGDepth is based on previous self-supervised monocular depth estimation algorithms, such as Monodepth godard2017unsupervised and Manydepth watson2021temporal , with the goal of estimating depth from monocular video recorded by an unknown camera.
-However, existing methods face difficulties in accurately estimating depth for moving objects lee2021learning ; li2021unsupervised .
-We propose the following innovative strategies to achieve accurate monocular depth estimation in complex dynamic scenarios.
-(1) We make use of the optical flow information and prior depth information to construct a new static reference frame.
-(2) Based on the new static reference frame, the original reference frame, and the new target frame, we construct a motion-guided cost volume.
-(3) We introduce an attention-based depth net structure, specifically designed to proficiently amalgamate feature maps with varying resolutions derived from a high-resolution network.
-3.2 Framework
+To tackle this limitation, we introduce Manydepth2, a technique adept at managing dynamic foreground by integrating temporal data into the cost volume and implementing a motion-guided photometric loss function.
+II-B Self-supervised VDE for Dynamic Objects
+Owing to the unique attributes of dynamic objects, there has been a focused effort among researchers to handle these objects distinctively throughout both the training and inference stages within self-supervised vision-based depth estimation (VDE) techniques.
+In [14], a deliberate decision was made to exclude dynamic objects from the analysis.
+This exclusion was aimed at preventing any potential interference with the optimization process, thereby ensuring the precision and accuracy of the depth estimation results.
+On the other hand, in [15], a more nuanced approach was adopted.
+Dynamic objects were first identified and segmented from the scene.
+Subsequent to this segmentation, these objects were treated distinctively within the framework of the photometric loss function.
+Such a strategy facilitates the integration of dynamic objects into the depth learning mechanism, ensuring that they contribute meaningfully without merely being sidelined or excluded.
+Furthermore, recent research efforts, as seen in [16], have ventured into predicting motion information at the object level.
+This predictive capability is then harnessed to establish more refined constraints, thereby enhancing the efficacy of self-supervised depth learning methodologies.
+Despite these advancements, it’s noteworthy that many of these approaches either grapple with intensive computational requirements or demonstrate only marginal enhancements in performance metrics.
+Recognizing these challenges, our novel approach, termed Manydepth2, seeks to overcome these constraints.
+We introduce a motion-guided cost volume framework that seamlessly integrates forecasted optical flow insights, all achieved through a sophisticated attention mechanism.
+III Methodology
+III-A Overview
+Firstly, as a foundational step, we utilize the potential of both optical flow information and pre-training depth data.
+By combining these insights, we construct a novel static reference frame.
+This innovative approach ensures that the depth estimation process remains robust and relevant, even when confronted with dynamic elements within the scene.
+With this static reference frame, our methodology further integrates the novel static reference frame alongside the original target frame.
+Combining these frames leads to creating a carefully designed motion-guided cost volume.
+This volume is crucial, improving the depth estimation by considering the detailed movement and space relationships in the scene.
+Finally, We are introducing a depth net structure that utilizes attention mechanisms.
+This structure is meticulously crafted to effectively combine feature maps that come from different resolutions, all originating from a high-resolution network.
+III-B Framework
 Our objective is to estimate the depth map ${D_{t}}$ and the rigid transformation ${[R|\bm{t}]_{rt}}$ between the target frame ${I_{t}}$ and a reference frame ${I_{r}}$.
@@
 This loss function serves as the optimization objective for the proposed neural network.
-The structure of MGDpth is demonstrated in Fig.
+The structure of Manydepth2 is demonstrated in Fig.
 2.
-At the outset, MGDpth employs a pretrained optical flow network denoted by $\theta_{f}$ and a pose network denoted by $\theta_{p}$ to predict the optical flow $f_{r\leftarrow t}$ and the transformation matrix $[R|\bm{t}]_{rt}$ from the target frame to reference frame.
+At the outset, Manydepth2 employs a trained optical flow network denoted by $\theta_{f}$ and a pose network denoted by $\theta_{p}$ to predict the optical flow $f_{r\leftarrow t}$ and the transformation matrix $[R|\bm{t}]_{rt}$ from the target frame to reference frame.
 Concurrently, a prior depth network denoted by $\theta_{cd}$ is utilized to estimate the coarse depth map $D_{t}^{c}$ for the target frame.
@@
 $p_{t}$,
-3.3 New Static Reference Frame
+III-C New Static Reference Frame (Rewrite and Check)
 To address the depth estimation of moving objects with monocular video, we leverage the estimation of flow $f_{r\leftarrow t}$, transformation matrix $[R|\bm{t}]_{rt}$, and coarse depth $D^{c}_{t}$.
@@
 Then the dynamic flow can be calculated as $f^{d}_{r\leftarrow t}=f_{r\leftarrow t}-f^{\prime}_{r\leftarrow t}$.
-The motion mask could be generated as follow:
+Figure 4: Qualitative results on KITTI Dataset.
+The initial row displays the RGB images of target frames where depth information has been estimated.
+The subsequent rows, specifically the second, third, and fourth, showcase depth maps generated by Monodepth2 trained with stereo techniques, ManyDepth, and Manydepth2, respectively.
+The motion mask could be generated as following:
 $\mathcal{M}_{m}^{t}=||f_{r\leftarrow t}-f^{\prime}_{r\leftarrow t}||^{2}_{2}>\epsilon,$ (3)
 where $\epsilon$ is the threshold for distinguishing moving parts.
-The motion pattern observed in the reference frame can be utilized to create a corresponding mask in the target frame.
-This is achieved through the generation of a static motion mask $\mathcal{M}_{ms}^{r}$, utilizing the static optical flow $f^{\prime}_{r\leftarrow t}$.
-Simultaneously, the dynamic motion mask $\mathcal{M}_{md}^{r}$ is derived with the assistance of $f_{r\leftarrow t}$.
-The composite motion mask for the reference frame is obtained by combining these two masks, denoted as $\mathcal{M}_{m}^{r}=\mathcal{M}_{ms}^{r}\cup\mathcal{M}_{md}^{r}$.
-The rationale for avoiding the direct computation of mask $\mathcal{M}_{m}^{r}$ using $f^{\prime}_{t\leftarrow r}$ lies in the necessity of inferring the depth of the target frame $D_{r}$.
-This procedure would subsequently lead to a rise in computational demands for both the training and inference processes.
-With $\mathcal{M}_{m}$, it becomes possible to reconstruct a new static reference image $\hat{I_{r}}$ by eliminating the impact of dynamic objects. This process can be formulated as follows:
-$\hat{I_{r}}=\begin{cases}I_{r}&p_{i}\notin\mathcal{M}_{m};\\ \phi^{\prime}(I_{t},D_{t},K_{t},K_{r},[R|\bm{t}]_{rt})&p_{i}\in\mathcal{M}_{m}.\\ \end{cases}$ (4)
+The motion mask in the target mask can be used to generate the motion mask $\mathcal{M}m^{r}$ in the reference mask, utilizing $f{r\leftarrow t}$.
+With $\mathcal{M}_{m}^{r}$, it becomes possible to reconstruct a new static reference image $\hat{I_{r}}$ by eliminating the impact of dynamic objects.
+This process can be formulated as follows:
+$\hat{I_{r}}=\begin{cases}I_{r}&p_{i}\notin\mathcal{M}_{m}^{r};\\ \phi^{\prime}(I_{t},D_{t},K_{t},K_{r},[R|\bm{t}]_{rt})&p_{i}\in\mathcal{M}_{m}^{r}.\\ \end{cases}$ (4)
 It is important to note that $\phi^{\prime}(.)$ differs from $\phi(.)$, as $\phi(.)$ is utilized to generate current frame information via sampling based on pixel correspondence between the current frame and the adjacent frame, whereas $\phi^{\prime}(.)$ generates adjacent frame by shifting forward pixels of the current frame.
-Please note that $\hat{I_{r}}$ is detached to eliminate its impact on the gradient calculation of backpropagation process.
-3.4 Motion-Guided Cost Volume
-Figure 4: Illustration for the construction of Final Volume.
-Feature maps of the target, reference, and new static reference frames $\mathcal{F}_{t}$, ${\mathcal{F}_{r}}$ and $\hat{\mathcal{F}_{r}}$ are used to generate motion-guided cost volume $\mathcal{V}_{m}$ with the help of $\mathcal{D}$ and $[R|t]_{rt}$.
-This volume $\mathcal{V}_{m}$ is combined with target feature map $\mathcal{F}_{t}$ via a channel attention to form the Final Volume.
+III-D Motion-Guided Cost Volume
 In constructing the motion-guided cost volume $\mathcal{V}_{m}$ for target frame $I_{t}$, a set of parallel planes that are perpendicular to the optic axis of $I_{t}$ is defined, based on the depth assumptions $\mathcal{D}=\{d_{k};k=1,...,M\}$, where $M$ represents the number of planes.
-The feature extractor $\theta_{fe}$ is utilized to generate the feature maps $\mathcal{F}_{t}$, $\mathcal{F}_{r}$ and $\hat{\mathcal{F}_{r}}$ of $I_{t}$, $I_{r}$ and $\hat{I_{r}}$, respectively.
-With the aid of $\mathcal{D}$ and $[R|\bm{t}]_{rt}$, a set of wrapped feature maps $\{\hat{\mathcal{F}_{t\leftarrow r}^{d_{k}}},\mathcal{F}_{t\leftarrow r}^{d_{k}};d_{k}\in\mathcal{D}\}$ are generated by warping $\hat{\mathcal{F}_{r}}$ and $\mathcal{F}_{r}$.
+The feature extractor $\theta_{fe}$ is utilized to generate the feature maps $\mathcal{F}_{t}$ and $\hat{\mathcal{F}_{r}}$ of $I_{t}$ and $\alpha\hat{I_{r}}+(1-\alpha)I_{r}$, respectively.
+With the aid of $\mathcal{D}$ and $[R|\bm{t}]_{rt}$, a set of wrapped feature maps $\{\hat{\mathcal{F}_{t\leftarrow r}^{d_{k}}};d_{k}\in\mathcal{D}\}$ are generated by warping $\hat{\mathcal{F}_{r}}$.
 Considering $\mathcal{V}_{m}=\{\mathcal{P}_{k},k=1,...,M\}$, it could be written as:
-$\mathcal{P}_{k}=\alpha\frac{\sum_{i=1}^{N}|\hat{\mathcal{F}_{t\leftarrow r}^{d_{k}}}-\mathcal{F}_{t}|}{N}+(1-\alpha)\frac{\sum_{i=1}^{N}|\mathcal{F}_{t\leftarrow r}^{d_{k}}-\mathcal{F}_{t}|}{N}.$ (5)
-Figure 5: Illustration for the construction of Final Volume.
-Feature maps of the target, reference, and new static reference frames $\mathcal{F}_{t}$, ${\mathcal{F}_{r}}$ and $\hat{\mathcal{F}_{r}}$ are used to generate motion-guided cost volume $\mathcal{V}_{m}$ with the help of $\mathcal{D}$ and $[R|t]_{rt}$.
-This volume $\mathcal{V}_{m}$ is combined with target feature map $\mathcal{F}_{t}$ via a channel attention to form the Final Volume.
-Here, $N$ denotes the number of reference frames used in construction, while in our experiment, we typically use a single adjacent frame as a reference frame.
+$\mathcal{P}_{k}=\frac{\sum_{i=1}^{N}|\hat{\mathcal{F}_{t\leftarrow r}^{d_{k}}}-\mathcal{F}_{t}|}{N}$ (5)
+Here, $N$ denotes the number of reference frames used in construction, while in our experiment.
 $\alpha$ denotes the parameters used to balance the influence of new static reference frame $\hat{I_{r}}$.
-The main reason for keeping the first item in the motion-guided cost volume is that at the early stage of the training process, the prior depth is not accurate enough which makes the $f^{\prime}_{r\leftarrow t}$ noisy.
-With the progress of training, $\alpha$ will gradually decrease to improve the influence of the new static reference frame in the construction of $\mathcal{V}_{m}$.
 Instead of directly taking the volume $\mathcal{V}_{m}$ into the later depth network, we propose to construct the final feature volume $\mathcal{V}_{f}$ by combining the $\mathcal{V}_{m}$ with the $\mathcal{F}_{t}$ though channel attention.
 The process of constructing the final feature volume $\mathcal{V}_{f}$ is presented in Fig.
-5.
-Figure 6: Illustration for the structure of Attention-based Depth Net.
-To effectively leverage feature maps of varying resolutions, we employ efficient channel attention to concatenate high-resolution feature maps with low-resolution feature maps and the low-resolution disparity map.
-This resultant feature map is subsequently utilized for disparity prediction within the current branch and is propagated to the subsequent branch.
-3.5 Attention-Based Depth Network
-The High-Resolution Network (HRNet) wang2020deep is well-regarded for its ability to preserve a high level of detail in input images.
-Although it is widely used in various computer vision tasks, its performance in Vision Depth Estimation (VDE) has not been extensively explored.
-To enhance the accuracy and efficiency of VDE, we integrate an efficient channel attention module into the design of the depth decoder network of the HRNet-based depth network.
+2.
+III-E Attention-Based Depth Network
+The High-Resolution Network (HRNet) [17] is well-regarded for its ability to preserve a high level of detail in input images.
 The HRNet is composed of multiple branches, denoted as $B$, each generating $S=b,...,4$ features with resolutions of $(\frac{H}{2^{b-1}},\frac{W}{2^{b-1}})$.
-Following the framework of Monodepth2, we utilize all branches to produce multi-scale disparity maps denoted as $D^{b}$, and apply the same loss function to each branch.
-However, instead of exclusively utilizing the feature map from the last stage of each branch for disparity map prediction, we leverage the efficient channel attention mechanism.
-This mechanism integrates feature maps from the current branch’s various stages, feature maps from deeper branches, and disparities from deeper branches.
-This fusion process can be expressed as $\mathcal{A}([f^{b}_{s}]_{s=b}^{s=4},[f^{b+1}_{s}]_{s=b+1}^{s=4},D^{b+1})$, as in Fig.
-6 for the $b^{th}$ branch, to collectively inform the disparity estimation process.
-3.6 Loss Function
-To summarize, there are three loss functions used in updating MGDepth’s weights. The final loss function $\mathcal{L}$ could be written as:
+However, instead of exclusively utilizing the feature map from the last stage of each branch for disparity map prediction, we leverage the channel attention mechanism.
+This mechanism integrates feature maps from the current branch’s various stages and feature maps from deeper branches.
+This fusion process can be expressed as $\mathcal{A}([f^{b}_{s}]_{s=b}^{s=4},[f^{b+1}_{s}]_{s=b+1}^{s=4})$.
+III-F Loss Function
+To summarize, there are three loss functions used in updating Manydepth2’s weights. The final loss function $\mathcal{L}$ could be written as:
 $\mathcal{L}=\mathcal{L}_{p}+\mathcal{L}_{s}+\mathcal{L}_{c}.$ (6)
@@
 $\begin{split}\mathcal{L}_{p}&=0.15*\sum_{p}\mathcal{M}_{o}\odot|I_{t}-I^{\prime}_{t}|\\ &+0.85*\frac{1-\text{SSIM}(I_{t},I^{\prime}_{t})}{2},\end{split}$ (7)
-where $\mathcal{M}_{o}$ is the auto mask introduced in godard2017unsupervised . $\mathcal{L}_{s}$ is the smooth loss, which could be written as:
+where $\mathcal{M}_{o}$ is the auto mask introduced in [18]. $\mathcal{L}_{s}$ is the smooth loss, which could be written as:
 $\mathcal{L}_{s}=|\partial_{x}D_{t}^{*}|e^{-|\partial_{x}I_{t}|}+|\partial_{y}D_{t}^{*}|e^{-|\partial_{y}I_{t}|},$ (8)
-where $d_{t}^{:}*=d_{t}/\overline{d_{t}}$ and $\overline{d_{t}}$ is the average depth of target frame wang2018learning .
-And $\mathcal{L}_{c}$ is the consistency loss introduced by watson2021temporal , to preserve the consistency between prior monocular depth and final multi-frame depth, which could be written as:
+where $d_{t}^{:}*=d_{t}/\overline{d_{t}}$ and $\overline{d_{t}}$ is the average depth of target frame [19].
+And $\mathcal{L}_{c}$ is the consistency loss introduced by [7], to preserve the consistency between prior monocular depth and final multi-frame depth, which could be written as:
 $\mathcal{L}_{c}=\sum_{p}\mathcal{M}_{m}^{t}\odot|D_{t}^{c}-D_{t}|.$ (9)
-3.7 Network Structure
-3.7.1 Depth Prior Network
-The architecture of the Depth Prior Network (DPN) encompasses a dual-component system consisting of a depth encoder and a depth decoder.
-The depth encoder is established employing the HRNet16 model wang2020deep , which has been pre-trained on the ImageNet dataset deng2009imagenet and functions as the foundational backbone.
-This encoder operates across four distinct branches, each corresponding to a distinct pyramidal scale.
-The feature representations within each of these scales are subsequently channeled to the depth decoder through a sequence of skip connections, notably influenced by the UNet framework ronneberger2015u .
-The depth decoder is characterized by an attention-based architecture as originally introduced in our primary manuscript.
-This component is designed to facilitate the fusion of encoded features derived from the depth encoder.
-3.7.2 Pose Network
-PoseNet exclusively comprises an encoder architecture utilizing a ResNet18 model he2016deep that has been pre-trained on the ImageNet dataset deng2009imagenet .
-Nonetheless, its distinctive attribute lies in its capacity to generate a 6-degree-of-freedom camera ego-motion vector, including a 3d translational vector and a 3d rotational vector.
-3.7.3 Depth Network
-In contrast to the Depth Prior Network, the depth network has a very similar layout.
-The key distinction is that, rather than analyzing just one frame at a time, it often processes two consecutive frames together.
-Once the first stage’s feature extraction is complete, these paired feature maps are employed to create a ”cost volume” as outlined in the main text.
-This cost volume is then blended with the feature representation of the target frame, resulting in a comprehensive ”final volume.” This final volume plays a crucial role in the subsequent process of building the depth map.
-3.7.4 Optical Flow Network
-The low-level computer vision community has favored the task of estimating the pixel-wise correspondence (Optic Flow) between two frames due to its applicability across various scenarios, without the need for specialized prior knowledge.
-Our paper involves adapting the pre-trained Gmflow xu2022gmflow on synthetic datasets, including the Flying Thing, Monkaa, and Driving datasets from the Scene Flow Dataset MIFDB16 and the Sintel dataset butler2012naturalistic .
-Notably, Gmflow used in MGDepth was not trained on any dataset used in our depth and ego-motion estimation experiments.
-The rationale behind adopting Gmflow lies in its lightweight nature and rapid inference capabilities, attributed to the inclusion of a dilatation correlation layer.
-This selection was made based on the consideration that integrating Gmflow into the existing depth estimation framework would not impose substantial computational overhead.
-Figure 7: Qualitative results on KITTI Dataset.
-The initial row displays the RGB images of target frames where depth information has been estimated.
-The subsequent rows, specifically the second, third, and fourth, showcase depth maps generated by Monodepth2 trained with stereo techniques, ManyDepth, and MGDepth, respectively.
-3.8 Discussion And Thinking
-MGDepth distinguishes itself from previous methods, such as feng2022disentangling , that utilize category information for segmentation and object tracking.
-Unlike those methods, MGDepth operates independently of any prior dataset knowledge.
-The optical flow module, pretrained solely on synthetic testing data, significantly enhances its ability to adapt to various scenarios, eliminating the need for extra labeling efforts.
-Furthermore, unlike typical motion segmentation modules relying on optical flow yang2021learning that demand substantial computational resources, MGDepth takes a different approach.
-We directly leverage the optical flow model and integrate necessary information, resulting in reduced computational demands.
-4 Experiment
-The primary experiments were conducted on the KITTI-2015 geiger2012we and Cityscapes silberman2012indoor datasets.
-(1) KITTI-2015: We use the Eigen split eigen2015predicting for both training and testing our model.
-To maintain consistency with prior studies zhou2017unsupervised ; zhou2021r , we utilized the same data selection approach and excluded all static images from the training set.
-This resulted in 39,810 training images and 4,424 test images.
-To ensure uniformity, we employed the camera matrix configuration described in godard2019digging , which applies a single intrinsic matrix to all images and sets the length focal to the average value.
-(2) Cityscapes: To ensure consistency with previous investigations watson2021temporal ; yang2018lego , we selected 69,731 images from the monocular sequences from Cityscapes cordts2016cityscapes as our training dataset, which were preprocessed into triples using scripts from zhou2017unsupervised .
-Our approach does not take into account stereo pairs or semantics.
-Performance assessment was based on the 1,525 test images, and the provided SGM hirschmuller2007stereo disparity maps were used.
-Similar to KITTI, we restricted depths to a max of 80m.
+IV Experiments
 T Method M S F W x H Abs Rel Sq Rel RMSE RMSE log $\delta<1.25$ $\delta<1.25^{2}$ $\delta<1.25^{3}$
-KITTI 2015
-Stereo
-EPC++luo2019every 1024 x 320 0.128 0.935 5.011 0.209 0.831 0.945 0.979
-FeatDepthshu2020feature 1024 x 320 0.099 0.697 4.427 0.184 0.889 0.963 0.982
-R-MSFM6zhou2021r 1024 x 320 0.108 0.753 4.469 0.185 0.888 0.963 0.982
-DFRzhan2018unsupervised 608 x 160 0.135 1.132 5.585 0.229 0.820 0.933 0.971
-Monocular
-Packnet-SFM guizilini20203d 1280 x 384 0.107 0.802 4.538 0.186 0.889 0.962 0.981
-Monodepth2 godard2019digging 1024 x 320 0.115 0.882 4.701 0.190 0.879 0.961 0.982
-Zhao zhao2020towards $\bullet$ 1024 x 320 0.113 0.704 4.581 0.184 0.871 0.961 0.984
-Chen chen2022compacted $\bullet$ $\bullet$ 1024 x 320 0.112 0.859 4.798 0.193 0.879 0.960 0.981
-FeatDepth shu2020feature 1024 x 320 0.104 0.729 4.481 0.179 0.893 0.965 0.984
-Wang wang2020self $\bullet$ 1024 x 320 0.106 0.773 4.491 0.185 0.890 0.962 0.982
-DevNet zhou2022devnet 1024 x 320 0.103 0.713 4.459 0.177 0.890 0.965 0.982
-Ranjan ranjan2019competitive $\bullet$ 832 x 256 0.148 1.149 5.464 0.226 0.815 0.935 0.973
-EPC++ luo2019every 832 x 256 0.141 1.029 5.350 0.216 0.816 0.941 0.976
-Guizilini guizilini2020semantically $\bullet$ 640 x 192 0.102 0.698 4.381 0.178 0.896 0.964 0.984
-Johnston johnston2020self 640 x 192 0.106 0.861 4.699 0.185 0.889 0.962 0.982
-Monodepth2 godard2019digging 640 x 192 0.115 0.903 4.863 0.193 0.877 0.959 0.981
-Packnet-SFM guizilini20203d 640 x 192 0.111 0.785 4.601 0.189 0.878 0.960 0.982
-Patil patil2020don $\bullet$ $\bullet$ 640 x 192 0.111 0.821 4.650 0.187 0.883 0.961 0.982
-Wang wang2020self $\bullet$ 640 x 192 0.106 0.799 4.662 0.187 0.889 0.961 0.982
-Li li2023learning $\bullet$ $\bullet$ 640 x 192 0.102 0.703 4.348 0.175 0.895 0.966 0.984
-ManyDepthwatson2021temporal $\bullet$ 640 x 192 0.098 0.770 4.459 0.176 0.900 0.965 0.983
-Lee lee2023multi $\bullet$ 640 x 192 0.096 0.644 4.230 0.172 0.903 0.968 0.985
-DynamicDepth feng2022disentangling $\bullet$ $\bullet$ 640 x 192 0.096 0.720 4.458 0.175 0.897 0.964 0.984
-MGDepth $\bullet$ $\bullet$ 640 x 192 0.091 0.650 4.263 0.171 0.908 0.967 0.984
-Cityscapes
-Monocular
-Lee lee2021attentive 832 x 256 0.116 1.213 6.695 0.186 0.852 0.951 0.982
-InstaDM lee2021learning $\bullet$ 832 x 256 0.111 1.158 6.437 0.182 0.868 0.961 0.983
-Pilzer pilzer2018unsupervised 512 x 256 0.240 4.264 8.049 0.334 0.710 0.871 0.937
-Monodepth2 godard2019digging 416 x 128 0.129 1.569 6.876 0.187 0.849 0.957 0.983
-Videos in the Wild gordon2019depth 416 x 128 0.127 1.330 6.960 0.195 0.830 0.947 0.981
-Li li2021unsupervised 416 x 128 0.119 1.290 6.980 0.190 0.846 0.952 0.982
-Struct2Depth casser2019depth $\bullet$ 416 x 128 0.151 2.492 7.024 0.202 0.826 0.937 0.972
-ManyDepthwatson2021temporal $\bullet$ 416 x 128 0.114 1.193 6.223 0.170 0.875 0.967 0.989
-DynamicDepth feng2022disentangling $\bullet$ $\bullet$ 416 x 128 0.103 1.000 5.867 0.157 0.895 0.974 0.991
-MGDepth $\bullet$ $\bullet$ 416 x 128 0.097 0.792 5.827 0.154 0.903 0.975 0.993
-Table 1: Self-supervised depth estimation results on the KITTI 2015 and Cityscapes.
+KITTI 2015 Stereo EPC++[20] 1024 x 320 0.128 0.935 5.011 0.209 0.831 0.945 0.979
+FeatDepth[21] 1024 x 320 0.099 0.697 4.427 0.184 0.889 0.963 0.982
+DFR[22] 608 x 160 0.135 1.132 5.585 0.229 0.820 0.933 0.971
+Monocular Packnet-SFM [23] 1280 x 384 0.107 0.802 4.538 0.186 0.889 0.962 0.981
+Monodepth2 [24] 1024 x 320 0.115 0.882 4.701 0.190 0.879 0.961 0.982
+FeatDepth [21] 1024 x 320 0.104 0.729 4.481 0.179 0.893 0.965 0.984
+DevNet [9] 1024 x 320 0.103 0.713 4.459 0.177 0.890 0.965 0.982
+Ranjan [25] $\bullet$ 832 x 256 0.148 1.149 5.464 0.226 0.815 0.935 0.973
+EPC++ [20] 832 x 256 0.141 1.029 5.350 0.216 0.816 0.941 0.976
+Guizilini [26] $\bullet$ 640 x 192 0.102 0.698 4.381 0.178 0.896 0.964 0.984
+Monodepth2 [24] 640 x 192 0.115 0.903 4.863 0.193 0.877 0.959 0.981
+Packnet-SFM [23] 640 x 192 0.111 0.785 4.601 0.189 0.878 0.960 0.982
+Patil [11] $\bullet$ $\bullet$ 640 x 192 0.111 0.821 4.650 0.187 0.883 0.961 0.982
+Li [27] $\bullet$ $\bullet$ 640 x 192 0.102 0.703 4.348 0.175 0.895 0.966 0.984
+ManyDepth[7] $\bullet$ 640 x 192 0.098 0.770 4.459 0.176 0.900 0.965 0.983
+Lee[28] $\bullet$ 640 x 192 0.096 0.644 4.230 0.172 0.903 0.968 0.985
+AQUANet[29] $\bullet$ 640 x 192 0.105 0.621 4.227 0.179 0.889 0.964 0.984
+DCPI-Depth[30] $\bullet$ $\bullet$ 640 x 192 0.095 0.662 4.274 0.170 0.902 0.967 0.985
+DynamicDepth [15] $\bullet$ $\bullet$ 640 x 192 0.096 0.720 4.458 0.175 0.897 0.964 0.984
+Manydepth2 $\bullet$ $\bullet$ 640 x 192 0.091 0.649 4.232 0.170 0.909 0.968 0.984
+TABLE I: Self-supervised depth estimation results on the KITTI 2015.
 We evaluate the performance of algorithms utilizing either stereo or monocular information during the training phase.
 Metrics denoted by Color have a lower optimal value, while metrics denoted by Color have a higher optimal value.
-The upper portion shows results on the KITTI 2015 dataset and the lower portion shows results on the Cityscapes dataset.
 The optimal results in each subsection are denoted in bold.
@@
 F: Optical Flow.
-(3) KITTI Odometry: For the purpose of odometry evaluation, we subject our system to the official KITTI odometry split, encompassing 11 driving sequences.
-These sequences are accompanied by ground truth odometry acquired through IMU and GPS readings.
-In line with prior research endeavors zhou2017unsupervised ; zhou2021r , our model’s training is conducted on sequences 00-08, while sequences 09-10 are reserved for testing.
-The primary focus of this section is to evaluate the ability of MGDepth and other state-of-the-art methods to predict depth information and estimate ego-motion.
-The depth prediction results are evaluated using widely used metrics godard2019digging ; eigen2014depth .
-The initial evaluation involves assessing MGDepth’s depth estimation performance on the KITTI-2015 and Cityscapes.
-Next, we evaluate its ability to estimate depth for dynamic objects on Cityscapes, utilizing the ground truth motion segmentation map.
-Following this, we conduct ablation studies to investigate the effectiveness of each module proposed in our paper.
-Finally, we present the results of MGDepth’s ego-motion estimation performance on the KITTI Odometry dataset.
-4.1 Monocular Depth Estimation
-KITTI-2015: In this experiment, the ResNet18 he2016deep architecture was utilized as the backbone for the pose network.
-The HRNet16 was used as the backbone for coarse depth network and attention-based depth network.
-During the training phase, our MGDepth receives input images with a resolution of $192\times 640$.
-For testing, we set the minimum and maximum depth values to 0.1m and 80m, respectively, following the recommendations of godard2019digging ; jin2001real .
-In our experiment, the scale factor is computed using the median value of the ground-truth image, employing the same approach as in zhou2017unsupervised .
-The upper section of Tab.
-1 presents a comparison between the performance of MGDepth and state-of-the-art algorithms.
-Overall, MGDepth exhibits superior performance compared to other methods, irrespective of whether monocular or multiview images are used.
-In particular, MGDepth demonstrates a significant performance advantage over Manydepth, which also employs multiview images and a cost volume structure to estimate depth.
-MGDepth outperforms Manydepth by a substantial margin of $7.2\%$ in terms of Abs Rel.
-In comparison to Dynamicdepth, which employs semantic motion segmentation maps to leverage high-level computer vision information for precise depth estimation, MGDepth achieves a superior performance of $5.3\%$ in terms of the Abs Rel, without the inclusion of any domain-specific information.
+T Method M S F W x H Abs Rel Sq Rel RMSE RMSE log $\delta<1.25$ $\delta<1.25^{2}$ $\delta<1.25^{3}$
+Cityscapes Monocular Lee [31] 832 x 256 0.116 1.213 6.695 0.186 0.852 0.951 0.982
+InstaDM [16] $\bullet$ 832 x 256 0.111 1.158 6.437 0.182 0.868 0.961 0.983
+Pilzer [32] 512 x 256 0.240 4.264 8.049 0.334 0.710 0.871 0.937
+Monodepth2 [24] 416 x 128 0.129 1.569 6.876 0.187 0.849 0.957 0.983
+Videos in the Wild [33] 416 x 128 0.127 1.330 6.960 0.195 0.830 0.947 0.981
+Li [34] 416 x 128 0.119 1.290 6.980 0.190 0.846 0.952 0.982
+Struct2Depth [10] $\bullet$ 416 x 128 0.151 2.492 7.024 0.202 0.826 0.937 0.972
+ManyDepth[7] $\bullet$ 416 x 128 0.114 1.193 6.223 0.170 0.875 0.967 0.989
+DynamicDepth [15] $\bullet$ $\bullet$ 416 x 128 0.103 1.000 5.867 0.157 0.895 0.974 0.991
+Manydepth2 $\bullet$ $\bullet$ 416 x 128 0.097 0.792 5.827 0.154 0.903 0.975 0.993
+TABLE II: Self-supervised depth estimation results on the Cityscapes.
+We evaluate the performance of algorithms utilizing either stereo or monocular information during the training phase.
+Metrics denoted by Color have a lower optimal value, while metrics denoted by Color have a higher optimal value.
+The optimal results in each subsection are denoted in bold.
+Our method outperforms all prior approaches in most metrics across all subsections, irrespective of whether the baselines utilize multiple frames during testing.
+M: Multi-frames.
+S: Motion Segmentation.
+F: Optical Flow.
+The main tests and evaluations for our study were carried out using two primary datasets: KITTI-2015 [35] and Cityscapes [36].
+In our experiments, the ResNet18 [37] architecture was utilized for the pose network.
+The HRNet16 was used as the backbone for the attention-based depth network.
+The pre-trained Gmflow [38] on the Sintel dataset [39] are used as optic flow network.
+For testing, we set the minimum and maximum depth values to 0.1m and 80m, as in [24, 40].
+In our experiment, the scale factor is computed using the median value of the ground-truth image as in [8].
+IV-A Monocular Depth Estimation
+KITTI-2015: The Tab.
+I presents a comparison between the performance of Manydepth2 and state-of-the-art algorithms.
+Overall, Manydepth2 exhibits superior performance compared to other methods, irrespective of whether monocular or multiview images are used.
+In particular, Manydepth2 demonstrates a significant performance advantage over Manydepth, which also employs multiview images and a cost volume structure to estimate depth.
+Manydepth2 outperforms Manydepth by a substantial margin of $7.2\%$ in terms of Abs Rel.
+In comparison to Dynamicdepth, which employs semantic motion segmentation maps to leverage high-level computer vision information for precise depth estimation, Manydepth2 achieves a superior performance of $5.3\%$ in terms of the Abs Rel, without the inclusion of any domain-specific information.
 Quantitative results are shown in Fig.
-7.
+4.
 Cityscapes: In contrast to the KITTI dataset, the Cityscapes dataset features a higher percentage of dynamic objects.
-During the training phase on the Cityscapes dataset, the input resolution of MGDepth is set to $128\times 416$.
+During the training phase on the Cityscapes dataset, the input resolution of Manydepth2 is set to $128\times 416$.
 The lower section of Tab.
-1 indicates a substantial performance advantage of MGDepth over its competing methods.
-Remarkably, MGDepth exhibits superior performance over ManyDepth, with a margin of approximately $15.0\%$ in the absolute relative error metric.
+II indicates a substantial performance advantage of Manydepth2 over its competing methods.
+Remarkably, Manydepth2 exhibits superior performance over ManyDepth, with a margin of approximately $15.0\%$ in the absolute relative error metric.
 Through the visualization of error maps generated by computing the differences between predicted depth maps and their corresponding ground truth depth maps, Fig.
 1 facilitates a qualitative analysis of depth predictions.
-These observations indicate that MGDepth offers enhanced accuracy for depth estimation of both dynamic foreground and static background.
-Additionally, the incorporation of a pre-trained optical flow estimation model with low computational requirements has resulted in a minimal increase in parameter size and running time for MGDepth.
+These observations indicate that Manydepth2 offers enhanced accuracy for depth estimation of both dynamic foreground and static background.
+Additionally, the incorporation of a pre-trained optical flow estimation model with low computational requirements has resulted in a minimal increase in parameter size and running time for Manydepth2.
 Quantitative results are shown in Fig.
-8.
-Type Method M WxH Abs Rel Sq Rel RMSE RMSE log $\delta<1.25$ $\delta<1.25^{2}$ $\delta<1.25^{3}$
-S InstaDM lee2021learning 832 x 256 0.139 1.698 5.760 0.181 0.859 0.959 0.982
-w/o S Monodepth2 godard2019digging 416 x 128 0.159 1.937 6.363 0.201 0.816 0.950 0.981
-ManyDepth watson2021temporal $\bullet$ 416 x 128 0.169 2.175 6.634 0.218 0.789 0.921 0.969
-MGDepth $\bullet$ 416 x 128 0.123 1.260 4.519 0.144 0.869 0.971 0.989
-Table 2: Results of Dynamic Foreground on Cityscapes.
-We evaluate depth prediction errors for dynamic objects such as vehicles, people, and bikes on the Cityscapes dataset silberman2012indoor .
+1.
+Tab.
+III showcases the depth errors within dynamic object regions by utilizing the ground truth motion segmentation maps provided in [16], which corresponds to the qualitative results presented in Fig.
+1.
+Type Method M WxH Abs Rel Sq Rel RMSE RMSE log
+S InstaDM [16] 832 x 256 0.139 1.698 5.760 0.181
+w/o S Monodepth2 [24] 416 x 128 0.159 1.937 6.363 0.201
+ManyDepth [7] $\bullet$ 416 x 128 0.169 2.175 6.634 0.218
+Manydepth2 $\bullet$ 416 x 128 0.123 1.260 4.519 0.144
+TABLE III: Results of Dynamic Foreground on Cityscapes.
+We evaluate depth prediction errors for dynamic objects such as vehicles, people, and bikes on the Cityscapes dataset [36].
 The best results are highlighted in bold.
-Figure 8: Qualitative results on Cityscapes (Error Map).
-The first row exhibits the RGB image of target frames for which depth information is estimated.
-The second and third rows illustrate error maps depicting the absolute relative error concerning the ground truth, where the degree of accuracy is represented by white (good) to red (poor) color gradations.
-The mask proposed in watson2021temporal is employed during drawing.
-4.2 Depth For Dynamic Foreground
+Method Tr R Tr R
+DFR [22] 11.93 3.91 12.45 3.46
+NeuralBundler [41] 8.10 2.81 12.90 3.71
+ManyDepth[7] 8.08 1.97 9.86 3.42
+Manydepth2 7.01 1.76 7.29 2.65
+TABLE IV: Visual odometry results on Seq. 9 and Seq. 10 of the KITTI odometry dataset. The average drift in root mean square error for both translation (Tr) and rotation (R) is reported.
+Ablation Abs Rel Sq Rel RMSE
+Final Prior Depth 0.105 0.768 4.536
+w/o Motion-Guided Volume 0.094 0.676 4.289
+w/o Attention for Final Volume 0.093 0.648 4.338
+w Normal HRNet 0.096 0.746 4.349
+Manydepth2 0.091 0.650 4.263
+TABLE V: Ablation study results of depth accuracy on the KITTI 2015. The proposed Manydepth2 model was evaluated by modifying its components.
+IV-B Odometry Estimation
+To evaluate odometry estimation results, we follow the split used in [8, 22].
+Specifically, we trained our model using Seq.
+00-08 from the KITTI odometry dataset and test methods on Seq.
+09-10.
 Tab.
-2 showcases the depth errors within dynamic object regions by utilizing the ground truth motion segmentation maps provided in lee2021learning , which corresponds to the qualitative results presented in Fig.
-8.
-Based on these results, it can be inferred that self-supervised monocular depth estimation methods from previous studies struggle to accurately estimate depth in dynamic scenes.
-This is likely due to their reliance on photometric loss, which assumes that the majority of foreground objects are static.
-In dynamic scenes, the photometric loss assumption is often violated, leading to difficulty in providing accurate depth estimates.
-However, MGDepth presents an effective solution to this issue by handling dynamic objects.
-Moreover, the motion-guided ego-motion loss function used in MGDepth enables accurate pose estimation, which aids in constructing cost volumes similar to stereo depth estimation.
-As a result, MGDepth achieves superior performance.
+IV shows the results, where the translational and rotational motion were evaluated by using the root mean square error (RMSE).
+Manydepth2, which uses less computational resources, outperforms other learning-based techniques that use more computational resources, such as FeatDepth.
+The results indicate that Manydepth2 achieves a significant reduction in RMSE by approximately $24.1\%$ for translational motion and $22.5\%$ for rotational motion, as compared to ManyDepth, on Seq.
+10.
+In support of this result, we present a visual depiction of the trajectory of both methods in Fig.
+5, which shows that Manydepth2’s trajectory exhibits considerably less drift than that of FeatDepth.
+The superior performance of Manydepth2 in precise pose estimation can be attributed to its use of both depth-based photometric loss and flow-based ego-motion loss, as well as the implementation of a motion-guided mask that effectively filters out dynamic foreground outliers.
+This combination of techniques enables Manydepth2 to achieve more reliable pose estimates compared to others.
 Figure 5: Trajectory visualization on Seq.
@@
 The ground truth trajectories are represented by red lines.
-Trajectories generated by the FeatDepth shu2020feature results are indicated by green lines, while our method’s trajectories are depicted in blue lines.
-Method Tr R Tr R
-DFR zhan2018unsupervised 11.93 3.91 12.45 3.46
-MonoDepth2 godard2019digging 10.85 2.86 11.60 5.72
-NeuralBundler li2019pose 8.10 2.81 12.90 3.71
-SC-SfMlearner zhou2017unsupervised 8.24 2.19 10.70 4.58
-FeatDepth shu2020feature 8.73$\dagger$ 2.11 10.65$\dagger$ 4.91
-ManyDepthwatson2021temporal 8.08 1.97 9.86 3.42
-MGDepth 7.01 1.76 7.29 2.65
-Table 3: Visual odometry results on Seq.
-9 and Seq.
-10 of the KITTI odometry dataset.
-The average drift in root mean square error for both translation (Tr) and rotation (R) is reported.
-$\dagger$ denotes the method that outperforms the results reported in the original paper.
-4.3 Odometry Estimation
-To evaluate odometry estimation results, we follow the split used in zhou2017unsupervised ; zhan2018unsupervised .
-Specifically, we trained our model using Seq.
-00-08 from the KITTI odometry dataset and test methods on Seq.
-09-10.
-Tab.
-3 shows the results, where the translational and rotational motion were evaluated by using the root mean square error (RMSE).
-MGDepth, which uses less computational resources, outperforms other learning-based techniques that use more computational resources, such as FeatDepth.
-The results indicate that MGDepth achieves a significant reduction in RMSE by approximately $24.1\%$ for translational motion and $22.5\%$ for rotational motion, as compared to ManyDepth, on Seq.
-10.
-In support of this result, we present a visual depiction of the trajectory of both methods in Fig.
-9, which shows that MGDepth’s trajectory exhibits considerably less drift than that of FeatDepth.
-The superior performance of MGDepth in precise pose estimation can be attributed to its use of both depth-based photometric loss and flow-based ego-motion loss, as well as the implementation of a motion-guided mask that effectively filters out dynamic foreground outliers.
-This combination of techniques enables MGDepth to achieve more reliable pose estimates compared to others.
-Ablation Abs Rel Sq Rel RMSE
-Final Prior Depth 0.108 0.753 4.563
-w/o Motion-Guided Volume 0.094 0.676 4.289
-w/o Attention for Final Volume 0.093 0.648 4.338
-w Normal HRNet 0.096 0.746 4.349
-MGDepth 0.091 0.650 4.263
-Table 4: Ablation study results of depth accuracy on the KITTI 2015. The proposed MGDepth model was evaluated by modifying its components.
-4.4 Abaltion Study
-We conduct ablation studies on the Cityscapes dataset to demonstrate the effectiveness of our proposed modules and validate our design choices.
-The results are presented in Tab.
-4, where the experiments were conducted under various settings, including Prior Depth Net of MGDepth, MGDepth without Motion-Guided Volume, MGDepth without Attention for Final Volume, and MGDepth with Normal HRNet.
-Tab.
-4 indicates that the addition of Attention-Based HRNet has a significant impact on the performance of MGDepth.
-And the introduction of Motion-Guided Volume and the Attention for Final Volume could further improve the performance of MGDepth.
-5 Conclusion
-We present MGDepth, a self-supervised multi-frame monocular depth prediction model leveraging optical flow-depth geometry for enhanced depth estimation.
-Our approach involves novel components such as a static reference frame, motion-guided cost volume, and an attention-based high-resolution neural network.
-MGDepth achieves accurate depth estimation for dynamic foreground and static background on KITTI-2015 and Cityscapes datasets.
-Notably, our results are reproducible using a single NVIDIA A10 GPU.
-MGDepth seamlessly integrates with existing self-supervised monocular depth estimation algorithms, serving as a refined coarse depth network.
-Moreover, our proposed modules augment accuracy by incorporating dynamic information.
-6 Acknowledgements
+Trajectories generated by the FeatDepth [21] results are indicated by green lines, while our method’s trajectories are depicted in blue lines.
+IV-C Abaltion Study
+In our research, we meticulously carry out ablation studies using the Cityscapes dataset.
+These studies are designed to dissect and showcase how effective our proposed modules are, allowing us to validate the specific design choices we’ve made.
+The detailed outcomes of these experiments are showcased in Table V.
+Within this table, we provide insights from experiments conducted under diverse conditions and configurations.
+Specifically, we assess Manydepth2’s performance with variations such as the Prior Depth Net, Manydepth2 configurations without Motion-Guided Volume, without Attention for Final Volume, and with the inclusion of a Normal HRNet.
+A key observation from Table V is the pronounced positive impact that integrating the Attention-Based HRNet has on enhancing Manydepth2’s performance.
+Furthermore, our findings underscore that the incorporation of both the Motion-Guided Volume and Attention for the Final Volume amplifies the capabilities of Manydepth2 even further, leading to improved outcomes.
+V Conclusion
+We introduce Manydepth2, an innovative self-supervised multi-frame monocular depth prediction model that harnesses the synergy of optical flow-depth geometry.
+One of the standout features of Manydepth2 is its unique components, including a static reference frame, motion-guided cost volume, and an attention-enhanced high-resolution neural network.
+When benchmarked against the KITTI-2015 and Cityscapes datasets, Manydepth2 consistently delivers precise depth estimation, effectively distinguishing between dynamic foreground elements and static backgrounds.
+Remarkably, our model’s performance is achieved using just a single NVIDIA A10 GPU, emphasizing its efficiency.
+Furthermore, Manydepth2 not only complements existing self-supervised monocular depth estimation techniques but also elevates them by integrating dynamic insights, thereby significantly enhancing accuracy.
+VI Acknowledgements
 Our research is supported by Amazon Web Services in the Oxford-Singapore Human-Machine Collaboration Programme and by the ACE-OPS project (EP/S030832/1).
-7 Declaration
-During the preparation of this work the authors used ChatGPT 3.5 in order to improve language and readability.
-After using this tool/service, the author(s) reviewed and edited the content as needed and take(s) full responsibility for the content of the publication.

### v3

@@
 where $\epsilon$ is the threshold for distinguishing moving parts.
-The motion mask in the target mask can be used to generate the motion mask $\mathcal{M}m^{r}$ in the reference mask, utilizing $f{r\leftarrow t}$.
+The motion mask in the target mask can be used to generate the motion mask $\mathcal{M}_{m}^{r}$ in the reference mask, utilizing $f_{r\leftarrow t}$.
 With $\mathcal{M}_{m}^{r}$, it becomes possible to reconstruct a new static reference image $\hat{I_{r}}$ by eliminating the impact of dynamic objects.
@@
 DynamicDepth [15] $\bullet$ $\bullet$ 640 x 192 0.096 0.720 4.458 0.175 0.897 0.964 0.984
+ManyDepth2-NF $\bullet$ 640 x 192 0.094 0.676 4.246 0.170 0.909 0.968 0.985
 Manydepth2 $\bullet$ $\bullet$ 640 x 192 0.091 0.649 4.232 0.170 0.909 0.968 0.984
@@
 F: Optical Flow.
+Manydepth2-NF refers to Manydepth2 without optical flow.
 T Method M S F W x H Abs Rel Sq Rel RMSE RMSE log $\delta<1.25$ $\delta<1.25^{2}$ $\delta<1.25^{3}$
@@
 4.
+We also present the results for Manydepth2-NF, a version of Manydepth2 that does not use optical flow or the motion-guided cost volume.
+We found that our attention-based network outperformed Manydepth by $4.1\%$.
 Cityscapes: In contrast to the KITTI dataset, the Cityscapes dataset features a higher percentage of dynamic objects.
@@
 Tab.
-III showcases the depth errors within dynamic object regions by utilizing the ground truth motion segmentation maps provided in [16], which corresponds to the qualitative results presented in Fig.
+III showcases the depth errors within dynamic object regions by utilizing the ground truth motion segmentation maps provided in [16], which corresponds to the qualitative results in Fig.
 1.
@@
 Final Prior Depth 0.105 0.768 4.536
-w/o Motion-Guided Volume 0.094 0.676 4.289
-w/o Attention for Final Volume 0.093 0.648 4.338
-w Normal HRNet 0.096 0.746 4.349
-Manydepth2 0.091 0.650 4.263
+w/o Attention for Volume 0.100 0.709 4.298
+Manydepth2-NF 0.094 0.676 4.246
+Manydepth2 0.091 0.649 4.232
 TABLE V: Ablation study results of depth accuracy on the KITTI 2015. The proposed Manydepth2 model was evaluated by modifying its components.
@@
 Within this table, we provide insights from experiments conducted under diverse conditions and configurations.
-Specifically, we assess Manydepth2’s performance with variations such as the Prior Depth Net, Manydepth2 configurations without Motion-Guided Volume, without Attention for Final Volume, and with the inclusion of a Normal HRNet.
+Specifically, we assess Manydepth2’s performance with variations such as the Prior Depth Net, Manydepth2 configurations without Motion-Guided Volume and without Attention for Final Volume.
 A key observation from Table V is the pronounced positive impact that integrating the Attention-Based HRNet has on enhancing Manydepth2’s performance.